### v1

# Continuous-Time Value Iteration for Multi-Agent Reinforcement Learning

Abstract
Existing reinforcement learning (RL) methods struggle with complex dynamical systems that demand interactions at high frequencies or irregular time intervals. Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differential value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation. While CTRL has shown promise, its applications have been largely limited to the single-agent domain. This limitation stems from two key challenges: (i) conventional solution methods for HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with HJB-based learning approaches, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training. In this paper, we propose a CT-MARL framework that uses physics-informed neural networks (PINNs) to approximate HJB-based value functions at scale. To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines $\nabla_{x}V$ along trajectories. This improves gradient fidelity, in turn yielding more accurate values and stronger policy learning. We evaluate our method using continuous-time variants of standard benchmarks, including multi-agent particle environment (MPE) and multi-agent MuJoCo. Our results demonstrate that our approach consistently outperforms existing continuous-time RL baselines and scales to complex multi-agent dynamics.
1 Introduction
RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation (Brunke et al., 2022), strategy games (Vinyals et al., 2019), wireless communications (Feriani and Hossain, 2021; Wang et al., 2023) , and traffic coordination (Haydari and Yılmaz, 2020). Most existing RL methods model these interactions in discrete time, where Bellman backup is computed at a fixed time interval. However, discrete-time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading (Shavandi and Khedmati, 2022) and autonomous driving (Kiran et al., 2021)). Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues (Wang et al., 2020; Shilova et al., 2024). To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time (Doya, 2000b; Rubanova et al., 2019). However, existing works only focus on single-agent settings and have not yet been widely explored for multi-agent scenarios. In multi-agent domains, each agent must not only interact with the environment but also coordinate with other agents, all while coping with the chllenges like non-stationarity introduced by simultaneously learning policies. Within this context, our paper investigates a novel approach to solving CT-MARL problems. To better understand the limitations of DTRL and advantages of CTRL for multi-agent scenarios, we present a didactic case as shown in Fig. 1. In this simple continuous-time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2. In contrast, our CT-MARL algorithm closely follows the ground-truth trajectory, maintains high returns, and generates accurate control actions for both agents.
(a) Value Approximation
(b) Reward Approximation
(c) Agent 1 Action
(d) Agent 2 Action
Figure 1: The performance of CT-MARL and DT-MARL is compared on a continuous-time, two-agent coupled oscillator task. DT-MARL suffers from significant bias and error when applied in the continuous domain. In contrast, CT-MARL yields smoother actions, higher rewards, and more accurate value approximations, closely aligning with the analytical LQR ground truth.
Unlike Bellman operator-based DTRL (Bellman, 1966), CTRL leverages HJB PDEs to compute differential value functions (Mukherjee and Liu, 2023; Shilova et al., 2024). However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems (Bellman et al., 1965), especially where the computational complexity grows exponentially with the state dimension in multi-agent systems. PINNs have emerged as a powerful tool to circumvent CoD (Weinan et al., 2021), and offer convergence guarantees for problems with smooth solutions (Shin et al., 2020; Ito et al., 2021). To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
However, even with PINNs to alleviate the CoD in CT-MARL, learning accurate centralized value functions under the centralized training decentralized execution (CTDE) framework remains challenging. The inherent non-stationarity across agents makes value learning unstable (Yu et al., 2022), and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients (Zhang et al., 2024b). These inaccuracies propagate into the value function, undermining policy quality and stability. To address this limitation, we propose a novel learning approach that combines PINN and VGI information, which is used to optimize the value learning. The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories. This integration yields significantly more accurate value-gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
Our work makes the following contributions. (1) We leverage PINNs to approximate differential value functions and apply them to solve continuous-time multi-agent reinforcement learning problems, which have rarely been explored by previous studies. (2) We introduce a novel value-gradient iteration term that dynamically refines the approximations of value gradients during training. This setting improves the computational accuracy of the value gradients, accelerates learning convergence, and leads to highly accurate value approximations, enabling efficient policy learning. (3) We create continuous-time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo. The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value-gradient learning in high-dimensional multi-agent systems.
2 Related Work
2.1 Continuous-Time Reinforcement Learning
CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings Jia and Zhou (2022); Lee and Sutton (2021); Wallace and Si (2023); Yang et al. (2021); Wiltzer et al. (2022); Vamvoudakis and Lewis (2010); Bian and Jiang (2021); Faradonbeh and Faradonbeh (2023). These studies aim to optimize policies in continuous-time domains, which avoids time discretization and achieves higher fidelity in control applications such as robotics and navigation. For instance, Bian and Jiang (2021) proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy. Similarly, Faradonbeh and Faradonbeh (2023) designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees. Lee and Sutton (2021) designs two policy iteration algorithms that compute comprehensive solutions to HJB equations. In contrast, Jia and Zhou (2022) introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization. Yang et al. (2021) develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics. Shilova et al. (2024) utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings. Yildiz et al. (2021) integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization. CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings. Luviano and Yu (2017) solves multi-agent pathfinding problem using fuzzy Q-learning, while Jiang et al. (2023) proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems. Beyond these examples, only a limited number of studies have addressed CT-MARL, which highlights the importance of our contributions.
2.2 Solving HJB Equations via PINNs
In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations Crandall and Lions (1983), which are the first-order nonlinear parabolic PDEs. However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD Osher et al. (2004); Osher and Shu (1991). Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth Weinan et al. (2021). PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual Han and Long (2020); Han et al. (2018), PDE residual Bansal and Tomlin (2021); Zhang et al. (2024b), and supervised data derived from numerical solvers Nakamura-Zimmerer et al. (2021); Zhang et al. (2024a). Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments Mukherjee and Liu (2023). However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
3 Methodology
3.1 Problem Formulation
In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework Doya (2000a); Yildiz et al. (2021); Lee and Sutton (2021), we formulate the continuous-time multi-agent problem as a tuple:
$\mathcal{M}=\bigl{\langle}\mathcal{X},\{\mathcal{U}_{i}\}_{i=1}^{N},N,f,r,g,\{t_{k}\}_{k\geq 0},T,\rho\bigr{\rangle}.$ (1)
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state space and $\mathcal{U}=\mathcal{U}_{1}\times\dots\times\mathcal{U}_{N}\subseteq\mathbb{R}^{m}$ represents the joint action space of $N$ agents. The global state and control input are represented by $x\in\mathcal{X}$ and $u\in\mathcal{U}$. Agent interactions occur over a fixed time horizon [0, T]. The multi-agent system evolves according to time-invariant nonlinear dynamics defined by $\dot{x}=f(x,u)$, where $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ is the global dynamics function. We define $\pi:\mathcal{X}\times[0,T]\rightarrow\mathcal{U}$ as the decentralized joint policy $\pi=(\pi_{1},\dots,\pi_{N})$. All agents share a global reward $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ and a global terminal cost $g:\mathcal{X}\rightarrow\mathbb{R}$. $\rho\in(0,1]$ is the discount factor or time-horizon scaling parameter. Unlike standard formulations that assume a fixed time step, we consider a strictly increasing sequence of decision times $\{t_{k}\}_{k\geq 0}$ with variable gaps $\tau_{k}=t_{k+1}-t_{k}>0$. In this paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f$ is Lipschitz continuous; $r$ and $g$ is Lipschitz continuous and bounded.
Definition 1 (Value Function of Multi-agent Systems).
Given $u=(u_{1},\ldots,u_{N})$ as a joint control input, the optimal global value function is defined as:
$V(x)=\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau)d\tau$ (2)
3.2 HJB and Policy Learning
For CT-MARL problems, we adopt the HJB theory to solve continuous-time problems. In this subsection, we explain how the HJB equation is leveraged to solve the CT-MARL problems. Specifically, we define a value network $V_{\theta}$ parametrized by a set of weights and biases $\theta$, and describe how the global value function $V_{\theta}$ is trained and how each agent’s policy $\pi_{\phi_{i}}$ is updated, such that the overall procedure serves as a continuous-time analogue of actor–critic policy iteration.
3.2.1 Critic Learning with HJB.
First, we present the following Lemma to show that the optimal value function $V(x,t)$, as defined in Eq. (2), is the optimal solution to HJB PDEs Crandall and Lions (1983) in the context of cooperative multi-agent settings.
Lemma 3.1 (HJB for Multi-agent Systems).
For all $(x,t)\in\mathcal{X}\times[0,T]$, the value function $V(x,t)$ is the optimal solution to satisfy the HJB PDEs and boundary conditions:
$\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f\bigl{(}x,u^{*}\bigr{)}+r\bigl{(}x,u^{*}\bigr{)}=0,$ (3)
where optimal control input $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,t,\nabla_{x}V(x,t))$, and boundary condition is $V(x,T)=g(x)$. The Hamiltonian $\mathcal{H}$ is defined as $\mathcal{H}=\nabla_{x}V(x,t)^{\!\top}\!f\bigl{(}x,u\bigr{)}+r\bigl{(}x,u\bigr{)}$.
The proof is attached in Appendix A.1.
To approximate differential value functions, we aim to solve the HJ PDEs given in Eq. (3). However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 Bui et al. (2022). To address this challenge, we adopt PINNs, which approximate value functions to satisfy both PDE residuals and boundary conditions. Specifically, we define HJB PDE residual as
$\mathcal{R}_{\theta}(x,t)=-\rho V_{\theta}(x,t)+\partial_{t}V_{\theta}(x,t)+\nabla_{x}V_{\theta}(x,t)^{\!\top}f(x,u)+r(x,u).$ (4)
and minimize both residual loss and boundary condition loss towards zero during model refinement.
$\mathcal{L}_{\text{res}}=\bigl{\|}\mathcal{R}_{\theta}(x,t)\bigr{\|}_{1},\quad\mathcal{L}_{\text{term}}=\bigl{\|}V_{\theta}(x,T)-g(x)\bigr{\|}_{2}^{2}.$ (5)
3.2.2 Policy Learning
While analytical optimal control laws can be derived in some cases by maximizing the Hamiltonian, such closed-form solutions might not be available in complex multi-agent systems such as MPE or MuJoCo. To overcome this challenge, we use an actor network to generate control inputs, replacing the need for analytical controls in the critic network used to approximate value functions. The actor and critic networks are refined iteratively until convergence, enabling the actor network to approximate optimal control policies. During training, we compute a continuous-time advantage function derived by residual $\mathcal{R}_{\theta}(x,t)$. This advantage function is used for policy gradient, where each agent’s decentralized policy $\pi_{\phi_{i}}(u_{i}\mid x,t)$ is optimized to maximize long-term return.
Lemma 3.2 (Instantaneous Advantage).
Assume the one-step Q-function over a short interval $\delta t>0$ be
$Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl{(}x(t+\delta t),t+\delta t\bigr{)}.$ (6)
Then the instantaneous advantage satisfies
$A(x,t,u)=-\rho V(x,t)+\partial V_{t}(x,t)+\nabla_{x}V(x,t)^{\!\top}\,f(x,u)+r(x,u).$ (7)
The proof is attached at Appendix A.2.
With the critic’s instantaneous advantage:
$A_{\theta}(x,t,u)=-\rho V_{\theta}(x,t)+\partial_{t}V_{\theta}(x,t)+\nabla_{x}V_{\theta}(x,t)^{\!\top}f(x,u)+r(x,u),$ (8)
we update each agent’s policy network $\pi_{\phi_{i}}(u_{i}\mid x,t)$ in a decentralized fashion. For agent $i$, we minimize the negative expected advantage under the joint policy:
$\mathcal{L}_{p_{i}}=-A_{\theta}\bigl{(}x,t,u\bigr{)}\log\pi_{\phi_{i}}\bigl{(}u_{i}\mid x,t\bigr{)},$ (9)
Here, $u=(u_{i},u_{-i})$ denotes the joint action, where $u_{i}\sim\pi_{\phi_{i}}$ is sampled from agent $i$’s policy, and $u_{-i}$ represents the actions of all other agents, sampled as $u_{-i}\sim\pi_{\phi_{-i}}$.
Lemma 3.3 (Policy Improvement).
Let $\pi_{\text{old}}$ be the current joint policy and $\pi_{\text{new}}$ the updated policy after one gradient step on the actor loss $\mathcal{L}_{p}$ with sufficiently small step size. Then:
$Q^{\pi_{\text{new}}}(x,t,u)\geq Q^{\pi_{\text{old}}}(x,t,u).$ (10)
The proof can be found at Appendix A.3.
3.3 Value Gradient Iteration Module
The performance of continuous-time control policies depends critically on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x,t)$. Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs Zhang et al. (2024b); Mukherjee and Liu (2023); Eberhard et al. (2024). However, the gradients obtained from standard PINN training are often noisy or misaligned with true trajectory behavior. Although prior work Zhang et al. (2024b) improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings. To address this, we introduce the VGI module that iteratively refines the value gradient approximations during training.
Definition 2 (VGI Gradient Estimator).
Given a small time step $\Delta t$, the VGI estimator of the value gradient at $(x_{t},u_{t})$ is defined by
$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\!\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$ (11)
The VGI target in Eq. (11) can be interpreted as a one-step unrolling of the Bellman equation in the space of gradients. The first term captures the instantaneous contribution of the local reward gradient, while the second term propagates the downstream value information through the Jacobian of the system dynamics. This construction resembles a semi-discretized version of the value gradient flow and provides a practical surrogate for supervised gradient learning in the absence of ground-truth derivatives. The derivation process is posted at Appendix A.4.
Theorem 3.4 (Convergence of VGI).
Let $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be defined as
$G(\zeta)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\top}\zeta,$ (12)
and assume the dynamics $\|\nabla_{x}f(x_{t},u_{t})\|$ is bounded . Then $G$ is a contraction, and the sequence $\zeta^{(k+1)}=G(\zeta^{(k)})$ converges to a unique fixed point $\zeta^{*}\in\mathbb{R}^{d}$.
The proof is detailed at Appendix A.5
Rather than introducing a separate network to predict value gradients, we directly compute the automatic derivative of the shared PINN critic $V_{\theta}(x,t)$. This gradient is then trained to match the VGI-generated target defined in Eq. (11). Specifically, we minimize the mean squared error between the computed and target gradients:
$\mathcal{L}_{\text{vgi}}=\bigl{\|}\nabla_{x}V_{\theta}(x_{t},t)-\hat{g}_{t}\bigr{\|}_{2}^{2},$ (13)
where $\hat{g}_{t}=\nabla_{x}r_{\phi}(x_{t},u_{t},\Delta t)\Delta t+e^{-\rho\Delta t}\nabla_{x}f_{\psi}(x_{t},u_{t},\Delta t)^{\!\top}\nabla_{x}V_{\theta}(x_{t+\Delta t},t+\Delta t).$ Here, $r_{\phi}(x,u,\Delta t)$ denotes a reward models and $f_{\psi}(x,u,\Delta t)$ represents a dynamics model, where $\phi$ and $\psi$ are respective network parameters.
3.4 Implementation Details
While the previous sections introduced our continuous-time actor–critic framework and the VGI module for value-gradient consistency, several practical considerations are essential to make the overall method operational and effective.
3.4.1 Dynamics Model and Reward Model
In a continuous-time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice. Instead, we adopt a discrete-time model-based approach Sutton (1991); Kaiser et al. (2019); Hafner et al. (2019) that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
$\mathcal{L}_{\text{dyn}}=\bigl{\|}f_{\psi}(x_{t},u_{t},\Delta t)-x_{t+\Delta t}\bigr{\|}_{2}^{2}.$ (14)
After learning $f_{\psi}$, we recover the continuous-time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
Similarly, we fit a reward network $r_{\phi}(x_{t},u_{t},\Delta t)$ to the observed instantaneous reward $r_{t}$:
$\mathcal{L}_{\text{rew}}=\bigl{\|}r_{\phi}(x_{t},u_{t},\Delta t)-r_{t}\bigr{\|}_{2}^{2}.$ (15)
Both $f_{\psi}$ and $r_{\phi}$ are trained jointly, enabling us to compute the VGI module’s target.
3.4.2 Anchor Loss for Critic Network
In addition to the HJB residual and terminal-condition losses, we incorporate a TD-style anchor loss to improve both the stability and accuracy of value learning. While the residual loss enforces the correctness of the value gradient, it does not constrain the value of $V(x,t)$. Terminal-condition losses can provide such supervision, but they often rely on access to well-defined terminal targets, which may be unavailable in complex continuous control environments such as MuJoCo. In these cases, the anchor loss offers an additional source of value landscape, helping the critic produce reasonable value approximations even when terminal rewards are sparse, delayed, or difficult to specify. We define the one-step return as:
$R_{t}=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V_{\theta}\bigl{(}x_{t+\Delta t},\,t+\Delta t\bigr{)}.$ (16)
The anchor loss then enforces the value network to match these returns:
$\mathcal{L}_{\text{anchor}}=\bigl{\|}V_{\theta}(x_{t},t)-R_{t}\bigr{\|}_{2}^{2}.$ (17)
The overall critic objective combines all four losses:
$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{res}}}_{\text{HJB residual}}\;+\;\underbrace{\lambda_{\text{term}}\,\mathcal{L}_{\text{term}}}_{\text{terminal BC}}\;+\;\underbrace{\lambda_{\text{anchor}}\,\mathcal{L}_{\text{anchor}}}_{\text{TD anchor}}\;+\;\underbrace{\lambda_{g}\,\mathcal{L}_{\text{vgi}}}_{\text{VGI consistency}}.$ (18)
Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value-bootstrap anchoring, and gradient consistency. In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory. The detailed training process is listed in Appendix A.6.
Figure 2: Performance across continuous-time multi-agent MuJoCo settings. The y-axis shows the mean cumulative reward.
4 Experimental Results
We evaluate our Value Iteration via PINN (VIP) method on two continuous-time multi-agent benchmarks: MPE Lowe et al. (2017b) and multi-agent MuJoCo Todorov et al. (2012); Peng et al. (2021). In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations. This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP. Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous-time RL baselines in these environments? (2) VGI ablation: How much does the VGI module contribute to final performance and training stability? (3) PINN design choice: How does activation function choice and loss term weighting in critic network affect the performance of VIP? (4) Time discretization impact: How well do discrete-time and continuous-time methods perform under arbitrary or unfixed time intervals?
Figure 3: $V$ and $\nabla_{x}V$ contour using VIP w/ VGI and w/o VGI in $d_{1}$-$d_{2}$ frame.
Benchmarks. We evaluate our method against baselines across eight experimental settings. We extend the MPE framework to a continuous-time formulation by using a variable-step Euler integration scheme, where the time interval $\Delta t$ is sampled from a predefined range at each step. Experiments are conducted on the cooperative navigation and predator prey environments. Similarly, we adapt MuJoCo to continuous-time settings and evaluate on a suite of multi-agent locomotion tasks, including ant (2 $\times$ 4, 2 $\times$ 4d, 4 $\times$ 2), walker, swimmer, and cheetah. Further implementation details of these benchmarks are provided in Appendix B.2 and B.3. Lastly, we introduce a simple yet illustrative coupled oscillator environment to highlight the behavior of exact value functions, value gradients, and the relative performance of different methods under a controlled setting.
4.1 Baseline Methods
To evaluate our CT-MARL framework VIP, we compare against three continuous-time policy iteration baselines and include an ablated variant of our method without VGI: CT-MBRL (ODE) Yildiz et al. (2021): A continuous-time model-based RL approach that learns system dynamics via Bayesian neural ODEs. This method uses an actor-critic framework to approximate state-value functions and derive the continuous-time optimal policies.
Figure 4: Performance across continuous-time MPE settings. The y-axis shows the mean cumulative reward.
Differential Policy Iteration (DPI) Lee and Sutton (2021): A model-based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations. Integral Policy Iteration (IPI) Lee and Sutton (2021): A partially model-free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement. In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions. Ablation (w/o VGI): An ablated version of VIP without VGI, which isolates the efficacy of value gradient refinement.
4.2 Results Analysis
Model Performance. We evaluate all RL methods in MPE and MuJoCo environments using five random seeds and report the mean cumulative reward curves in Fig. 2 and 4. The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL. As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations Krishnapriyan et al. (2021); Zhang et al. (2024b).
Figure 5: VIP performance with ReLU and Tanh activation functions in MuJoCo and MPE settings.
To address this limitation, we incorporate the anchor and VGI loss terms in Eq. (18), which capture the landscape of value and its gradient so that PINNs converge to the true values. Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments. This observation is consistent with the conclusion in the previous studies Zhang et al. (2024a), which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
To further demonstrate the importance of VGI, we revisit the didactic example, generate 400 rollouts from sampled initial states, and compute the average value using models with and without VGI. As shown in Fig. 3, the value contour projected onto the $d_{1}$$d_{2}$ frame reveals that the model with VGI closely matches the ground truth in both structure and scale, while the model without VGI produces significantly biased approximations. These results confirm that adding VGI is necessary to improve the accuracy of value approximations. A detailed comparison of the corresponding value gradients is also illustrated in Fig. 3. Furthermore, we evaluate the sensitivity of each loss term for the critic network by measuring the minimum distance to prey in Fig. 6, which highlights the critical role of PINN in refining value and policy networks.
Figure 6: Ablation study of different loss terms in critic network for cooperative predator prey.
Choice of PINN Design. We first examine the impact of activation function choice when incorporating PINN-based losses into critic network refinement. Specifically, we train VIP using ReLU and Tanh activations in both MuJoCo and MPE environments and report the accumulated rewards in Fig. 5. The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks. This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies Bansal and Tomlin (2021); Zhang et al. (2024b). The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs. PINNs typically use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs. The PINN residuals, including PDEs and boundary conditions, are further optimized using gradient descent. Smooth activation functions like Tanh support stable and accurate gradient flow throughout training, enabling more effective value approximations.
Figure 7: VIP performance with different weight settings in critic losses.
In contrast, VIP with ReLU often encounters zero-gradient regions during backpropagation, which results in gradient explosion for deeper network architectures or degrades the learning of value functions due to insufficient nonlinearity. Therefore, the choice of a smooth activation function like Tanh is better suited for physics-informed learning, thereby ensuring more accurate approximations of the value functions.
We also investigate the impact of weight parameters for PINN loss terms during VIP training. Specifically, we evaluate three configurations in Eq. (18): 1) balanced weights for all loss terms; 2) a large weight for boundary condition while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms. As shown in Fig. 7, the best accumulated reward performance is achieved when all loss terms are properly balanced. Imbalanced weight settings yield stiffness in the training dynamics of PINNs Wang et al. (2021), which makes the dominant loss term (with the largest weights) converge faster than the others Wang et al. (2022).
Figure 8: CT-MARL and DT-MARL performance under different time intervals.
In our experiments, such an imbalance causes VIP only to satisfy the PDEs or boundary conditions during training, ultimately leading to poor value approximations. Lastly, we evaluate the robustness of VIP and MADDPG Lowe et al. (2017a) by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts. Fig. 8 illustrates that VIP maintains a nearly constant return across different time intervals, whereas MADDPG’s performance degrades significantly as the interval increases. This result highlights the advantage of VIP in continuous-time multi-agent scenarios.
5 Conclusion and Limitations
We propose a novel approach that integrates PINNs into the actor-critic framework to solve CT-MARL problems. Specifically, we approximate value functions using HJB-based PINNs and introduce VGI to improve value approximations, thus mitigating the adverse impact of inaccurate value approximations on policy learning. We validate our VIP across continuous-time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA. Furthermore, we investigate the importance of activation function choice and loss term weighting for VIP performance. In summary, our proposed VIP offers a promising approach to solve CT-MARL problems. However, we notice the limitation of VIP: the current work only focuses on cooperative CT-MARL tasks. Many real-world multi-agent scenarios involve competitive interactions, which have not yet been explored within CT-MARL frameworks. In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN Zhang et al. (2024b). As part of future work, we aim to extend VIP to explore competitive CT-MARL problems, particularly those involving state constraints.
Appendix A Mathematical Proof
A.1 Proof of Lemma 3.1
Proof.
Given all $(x,t)\in\mathcal{X}\times[0,T]$ and a small horizon $\Delta t>0$, we apply a Taylor expansion to the definition of $V(x,t)$ in Eq. (2) to derive the HJB PDE as follows:
$V(x,t)$ $=\max_{u\in\mathcal{U}}\int_{t}^{T}e^{-\rho(\tau-t)}r(x,u)d\tau+g(x_{T})$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+g(x_{T})$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t}^{T-\Delta t}e^{-\rho(s+\Delta t-t)}r(x(s+\Delta t),u(s+\Delta t))ds+g(x_{T})$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}(\max_{u\in\mathcal{U}}\int_{t}^{T-\Delta t}e^{-\rho(s-t)}r(x(s+\Delta t),u(s+\Delta t))ds+g(x_{T}))$
$\approx\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}(\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+g(x_{T}))$
$\approx\max_{u\in\mathcal{U}}r(x,u)\Delta t+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}V(x(t+\Delta t),t+\Delta t)$
$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+(1-\rho\Delta t+o(\Delta t))\max_{u\in\mathcal{U}}(V(x,t)+\partial_{t}V(x,t)\Delta t+\nabla_{x}V(x,t)^{\!\top}\!f(x,u)\Delta t+o(\Delta t))$
By canceling out $V(x,t)$ on both sides of the above equality, we obtain that
$-\rho V(x,t)\Delta t+\partial_{t}V(x,t)\Delta t+\max_{u\in\mathcal{U}}(\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u))\Delta t+o(\Delta t)=0$
Dividing by $\Delta t$ and letting $\Delta t\to 0$, we have that
$-\rho V(x,t)+\partial_{t}V(x,t)+\max_{u\in\mathcal{U}}(\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u))=0.$
Therefore, the $V(x,t)$ is the optimal solution to the following HJB PDEs:
$\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f(x,u^{*})+r(x,u^{*})=0,$
with boundary condition $V(x,T)=g(x)$.
Here optimal control input is $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,t,\nabla_{x}V(x,t))$, where $\mathcal{H}$ is the Hamiltonian defined as $\mathcal{H}=\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u)$. ∎
A.2 Proof of Lemma 3.2
Proof.
Recall the one-step $Q$-function over a short interval $\delta t>0$, where $u$ is the optimal control input.
$Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}\,V\bigl{(}x(t+\delta t),\,t+\delta t\bigr{)}.$
For small $\delta t$ we have the first-order Taylor expansion in both state and time:
$V\bigl{(}x(t+\delta t),\,t+\delta t\bigr{)}=V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f(x,u)\,\delta t+\partial_{t}V(x,t)\,\delta t+o(\delta t).$
Similarly, $e^{-\rho\delta t}=1-\rho\,\delta t+o(\delta t)$.
Plugging both expansions into $Q(x,t,u)$ gives
$Q(x,t,u)$ $=r(x,u)\,\delta t+\bigl{(}1-\rho\,\delta t+o(\delta t)\bigr{)}\Bigl{[}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)\,\delta t+\partial_{t}V(x,t)\,\delta t+o(\delta t)\Bigr{]}$
$=r(x,u)\,\delta t+V(x,t)+\bigl{[}\nabla_{x}V(x,t)^{\!\top}f(x,u)+\partial_{t}V(x,t)-\rho V(x,t)\bigr{]}\delta t+o(\delta t).$
Subtract $V(x,t)$ and discard the higher-order term:
$Q(x,t,u)-V(x,t)=[-\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)+r(x,u)]\delta t+o(\delta t).$
Dividing by $\delta t$ and letting $\delta t\to 0$ yields the instantaneous advantage density
$A(x,t,u)=\lim_{\delta t\to 0}\frac{Q(x,t,u)-V(x,t)}{\delta t}=-\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)+r(x,u).$
This completes the proof. ∎
A.3 Proof of Lemma 3.3
Policy Improvement via State-Action Value Function.
We consider the standard policy improvement step, where the new policy is obtained by maximizing the state-action value function:
$\pi_{\text{new}}(x,t)=\arg\max_{u\in\mathcal{U}}Q^{\pi_{\text{old}}}(x,t,u),$
where the one-step Q-function with a small $\delta t>0$ is defined as:
$Q(x,t,u)$ $=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl{(}x(t+\delta t),t+\delta t\bigr{)}$
$=r(x,u)\,\delta t+e^{-\rho\delta t}\mathbb{E}_{u^{\prime}\sim\pi(\cdot\mid x^{\prime})}\left[Q(x^{\prime},t+\delta t,u^{\prime})\right],$
Since we assume the goal state function stays invariant, we only define the $Q^{\pi}=\int_{0}^{T}e^{-\rho t}r_{t}dt$. Because the new policy $\pi^{\text{new}}$ yields equal or higher value in expectation:
$\mathbb{E}_{u\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x,t,u)\right]\geq\mathbb{E}_{u\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x,t,u)\right].$
Then we can have:
$\begin{aligned} Q^{\pi_{\text{old}}}&=r_{0}\,\delta t+e^{-\rho\delta t}(\mathbb{E}_{u_{1}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{1},t_{1},u_{1})\right])\\ &\leq r_{0}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{1}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{1},t_{1},u_{1})\right]\right)\\ &=r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{2},t_{2},u_{2})\right]\right)\\ &\leq r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{2},t_{2},u_{2})\right]\right)\\ &=r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho t_{2}}r_{2}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{3}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{3},t_{3},u_{3})\right]\right)\\ &\vdots\\ &\leq\int_{0}^{T}e^{-\rho t}r_{t}dt\\ &=Q^{\pi_{\text{new}}}.\end{aligned}.$
A.4 Derivation of Definition 2
Proof.
We consider the value function defined in Eq. (2) and follow the Proof of Lemma 3.1 to write out the dynamic programming principle of $V(x,t)$ as:
$V(x_{t},t)=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V(x_{t+\Delta t},t+\Delta t).$
where $u$ is the optimal control input. Taking the gradient with respect to $x$ on both sides using the chain rule:
$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$
which matches the estimator proposed in Definition 2. ∎
A.5 Proof of Theorem 3.4
Proof.
From VGI in definition 2, with an Euler step gives $x_{t+\Delta t}=x_{t}+f(x_{t},u_{t})\Delta t+o(\Delta t)$, we have the first-order-accurate VGI approximation:
$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\bigl{[}I+\nabla_{x}f(x_{t},u_{t})\Delta t\bigr{]}^{\!\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$
Rewrite VGI approximation as the affine map $\zeta=G(\zeta)$ with
$b=\nabla_{x}r(x_{t},u_{t})\Delta t,\;A=e^{-\rho\Delta t}\bigl{[}I+\nabla_{x}f(x_{t},u_{t})\Delta t\bigr{]}^{\!\top},\;\zeta=\nabla_{x}V(\cdot,\cdot).$
Assume the dynamics have a bounded Jacobian $\|\nabla_{x}f(x_{t},u_{t})\|\leq L_{f}$. Then
$\|A\|\leq e^{-\rho\Delta t}(1+L_{f}\Delta t)=\beta.$
Because we study high-frequency settings, $\Delta t$ is chosen sufficiently small, which makes $L_{f}\Delta t\to 0$, so that $\beta<1$. For any $\zeta_{1},\zeta_{2}\in\mathbb{R}^{d}$,
$\|G(\zeta_{1})-G(\zeta_{2})\|=\|A(\zeta_{1}-\zeta_{2})\|\leq\beta\|\zeta_{1}-\zeta_{2}\|,$
Hence, $G$ is a contraction. Banach’s fixed-point theorem guarantees a unique fixed point $\zeta^{\ast}$ and linear convergence $\|\zeta^{(k)}-\zeta^{\ast}\|\leq\beta^{\,k}\|\zeta^{(0)}-\zeta^{\ast}\|$. Therefore, the value-gradient iteration converges, completing the proof. ∎
A.6 Training Algorithm
We present the training algorithm for our proposed approach, Value Iteration via PINN (VIP), as follows:
Algorithm 1 Value Iteration via PINN (VIP)
1:Init: value net $V_{\theta}$, policy nets $\{\pi_{\omega_{i}}\}_{i=1}^{N}$, dynamics $\hat{f}_{\psi}$, reward $\hat{r}_{\phi}$
2:for $l=1,\dots,T$ do
3:  $\triangleright$ Collect one rollout:
4:  $x\leftarrow\mathrm{env.reset}()$
5:  for $k=1,\dots,K$ do
6:   sample decision time $\sim\mathcal{T} \triangleright tt$ is arbitrary time
7:   for each agent $i=1,\dots,N$ do
8:     $u_{i}\sim\pi_{\omega_{i}}(u_{i}\mid x,t)$
9:   end for
10:   set joint action $u=(u_{1},\dots,u_{N})$
11:   $(x^{\prime},r)\leftarrow\mathrm{env.step}(u)$
12:   append $(x,t,u,r,x^{\prime})$ to local rollout $\mathcal{R}$
13:   $x\leftarrow x^{\prime}$
14:  end for
15:  $\triangleright$ Dynamics and Reward Model learning on $\mathcal{R}$
16:  update $\psi,\phi$ as per the Equation (14) and (15).
17:  $\triangleright$ Critic update on $\mathcal{R}$
18:  compute all losses $\mathcal{L}_{\mathrm{res}},\mathcal{L}_{\mathrm{term}},\mathcal{L}_{\mathrm{anchor}},\mathcal{L}_{\mathrm{vgi}}$ by Equation (5), (17), (13).
19:  $\theta\leftarrow\theta-\alpha_{V}\nabla_{\theta}(\ldots)$
20:  $\triangleright$ Actor update for each agent
21:  for $i=1,\dots,N$ do
22:   compute $A(x,t,u)$ for all $(x,t,u)\in\mathcal{R}$
23:   $\omega_{i}\leftarrow\omega_{i}-\alpha_{\pi}\,\nabla_{\omega_{i}}\bigl{(}-\mathbb{E}_{(x,t,u)\in\mathcal{R}}[\,A(x,t,u)\,\log\pi_{\omega_{i}}(u_{i}\mid x,t)\,]\bigr{)}$ by Equation (9).
24:  end for
25:end for
Appendix B Environmental Settings
B.1 Coupled Oscillator
We evaluate on a two-agent coupled spring–damper system. Each agent $i\in\{1,2\}$ controls one mass in a pair of identical oscillators with linear coupling. The continuous-time dynamics are
$\begin{aligned} \dot{x}_{i}&=v_{i},\\ \dot{v}_{i}&=-\,k\,x_{i}\;-\;b\,v_{i}\;+\;u_{i},\end{aligned}\qquad i=1,2,$ (19)
where
  • $x_{i}$ and $v_{i}$ are the position and velocity of mass $i$;
  • $k=1.0$ is the spring constant, and $b=0.5$ is the damping coefficient;
  • $u_{i}\in[-u_{\max},u_{\max}]$ is the control force applied by agent $i$, with $u_{\max}=10$.
At each step the joint reward is
$r=-\Bigl{(}x_{1}^{2}+x_{2}^{2}\;+\;\lambda_{c}\,(x_{1}-x_{2})^{2}\;+\;\beta\,(u_{1}^{2}+u_{2}^{2})\Bigr{)},$
with coupling strength $\lambda_{c}=2.0$ and control penalty $\beta=0.01$. We normalize by a constant factor (here $1/10$) so that $r\in[-1,0]$.
For the coupled oscillator with linear dynamics
$\dot{x}=A\,x+B\,u,$
we can compute the exact infinite-horizon LQR solution:
  1. 1.
    Solve the continuous algebraic Riccati equation (CARE)
    $A^{\top}P+P\,A-P\,B\,R^{-1}B^{\top}P+Q=0$
    for the symmetric matrix $P\in\mathcal{R}^{4\times 4}$.
  2. 2.
    Form the optimal state-feedback gain
    $K\;=\;R^{-1}B^{\top}P.$
  3. 3.
    The optimal control law is
    $u^{*}(x)\;=\;-\,K\,x,\qquad u_{i}^{*}=-K_{i}\,x,$
    where $K_{i}$ is the $i$-th row of $K$.
  4. 4.
    The corresponding optimal value function is the quadratic form
    $V^{*}(x)\;=\;x^{\top}P\,x,$
    whose exact gradient is
    $\nabla_{x}V^{*}(x)\;=\;2\,P\,x.$
We use $u^{*}(x)$, $V^{*}(x)$, and $\nabla_{x}V^{*}(x)$ as ground truth targets when evaluating the precision of the policy, the error of the value function, and the consistency of the gradient.
B.2 Continuous-Time MPE
We build on the standard MPE of Lowe et al. [2017b], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions. In the original MPE each control step advances the physics by a fixed time-step $\Delta t_{\rm fixed}=0.1\,$s. To evaluate our continuous-time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
$\Delta t_{k}\;\sim\;\mathrm{Uniform}(\Delta t_{\min},\,\Delta t_{\max}).$
The underlying dynamics, observation and action spaces, reward functions, and task definitions remain exactly as in the original MPE. For the cooperative predator-prey environment, we only control the predators (3 agents) action to capture the prey (1 agent). While the prey is set up with random actions.
B.3 Continuous-Time Multi-Agent MuJoCo
For high-dimensional control we adapt the discrete-time Multi-Agent MuJoCo suite (e.g. cooperative locomotion, quadruped rendezvous). By default MuJoCo uses an internal physics integrator with a base time-step of 0.010.01\,s and repeats each action for $K_{\rm fixed}=5$ frames, yielding an effective control interval $\Delta t_{\rm fixed}=0.05\,$s. We instead sample the number of frames per control step,
$K_{k}\;\sim\;\mathrm{UniformInteger}(1,\,9),$
so that each step advances by
$\Delta t_{k}=K_{k}\times 0.01\ \mathrm{s}\,\in[0.01,\,0.09]\mathrm{s}$
at random. All other aspects of the environment (observations, reward structure, termination conditions) are kept identical to the original multi-agent MuJoCo tasks.
Appendix C Additional Experimental Results
Experiments were conducted on hardware comprising an Intel(R) Xeon(R) Gold 6254 CPU @ 3.10GHz and four NVIDIA A5000 GPUs. This setup ensures the computational efficiency and precision required for the demanding simulations involved in multi-agent reinforcement learning and safety evaluations.
C.1 Value Gradient Comparison
(a) Ground Truth Gradient
(b) VPI Gradient w/ VGI
(c) VPI Gradient w/o VGI
Figure 9: Value gradient comparison between using of VPI module.
To further demonstrate the effectiveness of VIP, we sample the same trajectory in the coupled oscillator environment and compute the value gradient from the analytical LQR solution, from VIP equipped with the VGI module (ours), and from VIP trained without VGI, respectively. Fig. 9 presents the resulting 3-D surfaces. The surface in panel (b) preserves the principal ridges and valleys of the ground truth, showing that the network recovers the correct geometric structure of $\nabla_{x}V$; its absolute error remains below $0.02$ across almost all timesteps and gradient dimensions. In contrast, the surface in panel (c) is noticeably distorted: several peaks are flattened, troughs are misplaced, and the absolute error frequently exceeds $0.08$. This comparison confirms that the VIP module is critical for aligning the learned gradients with the analytical solution, thereby reducing bias and stabilising the HJB residual.
C.2 ReLU vs Tanh at Ant $2\times 4d$ and Ant $4\times 2$.
Figure 10: VIP performance with ReLU and Tanh at Ant $2\times 4d$ and Ant $4\times 2$.
Fig. 10 compares VIP’s learning curves when the policy network uses ReLU or Tanh activations on Ant $2\times 4d$ and Ant $4\times 2$. Across both tasks the Tanh implementation converges faster and attains a higher plateau reward, whereas the ReLU version peaks earlier and then undergoes a mild performance decay. The observation aligns with the earlier Tanh-versus-ReLU analysis reported in the main paper Fig. 5: smoother activation functions mitigate gradient saturation and promote more stable policy updates. The additional evidence from the two Ant variants therefore reinforces our previous claim that Tanh is better suited for value–gradient propagation in VIP.
Appendix D Hyper-parameters
Table 1: Hyperparameter settings used in all experiments.
Parameter Value
Episode length 50
Replay buffer size $10^{4}$
Discount factor $\rho$ 0.95
Soft update rate $\tau$ 0.001
Actor learning rate 0.0001
Critic learning rate 0.001
Dynamics model learning rate 0.001
Reward model learning rate 0.001
Exploration steps 1000
Model save interval 1000
Random seed 111-120
As Table LABEL:tab:hp shows, the exploration steps are used to delay the decay of the exploration rate: during the first 1000 steps, the exploration schedule remains fixed to encourage initial exploration. The soft update rate $\tau$ controls the target network update in the critic and value estimation, where we adopt a target network with an exponential moving average to stabilize bootstrapped training. This technique helps suppress oscillations in value learning and leads to more accurate estimation of long-horizon returns.
Table 2: Summary of neural network architectures used in our framework.
Network Input Dimension Architecture and Activation
Value Network State + Time ($d+1$) FC(128) → FC(128) → FC(1), ReLU or Tanh
Dynamics Network State + Joint Action + Time ($d+na+1$) FC(128) → FC(128) → FC($d$), ReLU
Reward Network State + Joint Action + Time ($d+na+1$) FC(128) → FC(128) → FC(1), ReLU
PolicyNet Observation + Time + Time Interval ($o+2$) FC(128) → FC(128) → FC(64) → FC($a$), ReLU
Table 2 summarizes the architectures of the neural networks used in our VIP framework. All networks are implemented as FC layers with hidden size 128 unless otherwise specified. The value network takes the concatenation of the global state and time $(d+1)$ as input and outputs a scalar value. In our implementation, we use the Tanh activation function for the value network, as it provides smoother and more stable gradient propagation, which is critical for PINN-based value approximation. To validate this choice, we conducted an ablation study comparing Tanh and ReLU activations in the previous section.

### v2

@@
 This improves gradient fidelity, in turn yielding more accurate values and stronger policy learning.
-We evaluate our method using continuous-time variants of standard benchmarks, including multi-agent particle environment (MPE) and multi-agent MuJoCo.
-Our results demonstrate that our approach consistently outperforms existing continuous-time RL baselines and scales to complex multi-agent dynamics.
+We evaluate our method using continuous‑time variants of standard benchmarks, including multi‑agent particle environment (MPE) and multi‑agent MuJoCo.
+Our results demonstrate that our approach consistently outperforms existing continuous‑time RL baselines and scales to complex multi-agent dynamics.
 1 Introduction
-RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation (Brunke et al., 2022), strategy games (Vinyals et al., 2019), wireless communications (Feriani and Hossain, 2021; Wang et al., 2023) , and traffic coordination (Haydari and Yılmaz, 2020).
+RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation [5], strategy games [39], wireless communications [12, 44] , and traffic coordination [16].
 Most existing RL methods model these interactions in discrete time, where Bellman backup is computed at a fixed time interval.
-However, discrete-time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading (Shavandi and Khedmati, 2022) and autonomous driving (Kiran et al., 2021)).
-Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues (Wang et al., 2020; Shilova et al., 2024).
-To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time (Doya, 2000b; Rubanova et al., 2019).
-However, existing works only focus on single-agent settings and have not yet been widely explored for multi-agent scenarios.
+However, discrete‑time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading [33] and autonomous driving [21]).
+Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues [41, 34].
+To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time [9, 32].
+However, existing works only focus on single‑agent settings and have not yet been widely explored for multi-agent scenarios.
 In multi-agent domains, each agent must not only interact with the environment but also coordinate with other agents, all while coping with the chllenges like non-stationarity introduced by simultaneously learning policies.
@@
 1.
-In this simple continuous-time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2.
-In contrast, our CT-MARL algorithm closely follows the ground-truth trajectory, maintains high returns, and generates accurate control actions for both agents.
+In this simple continuous‑time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2.
+In contrast, our CT‑MARL algorithm closely follows the ground‑truth trajectory, maintains high returns, and generates accurate control actions for both agents.
 (a) Value Approximation
@@
 In contrast, CT-MARL yields smoother actions, higher rewards, and more accurate value approximations, closely aligning with the analytical LQR ground truth.
-Unlike Bellman operator-based DTRL (Bellman, 1966), CTRL leverages HJB PDEs to compute differential value functions (Mukherjee and Liu, 2023; Shilova et al., 2024).
-However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems (Bellman et al., 1965), especially where the computational complexity grows exponentially with the state dimension in multi-agent systems.
-PINNs have emerged as a powerful tool to circumvent CoD (Weinan et al., 2021), and offer convergence guarantees for problems with smooth solutions (Shin et al., 2020; Ito et al., 2021).
+Unlike Bellman operator-based DTRL [2], CTRL leverages HJB PDEs to compute differential value functions [27, 34].
+However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems [3], especially where the computational complexity grows exponentially with the state dimension in multi-agent systems.
+PINNs have emerged as a powerful tool to circumvent CoD [45], and offer convergence guarantees for problems with smooth solutions [35, 17].
 To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
 However, even with PINNs to alleviate the CoD in CT-MARL, learning accurate centralized value functions under the centralized training decentralized execution (CTDE) framework remains challenging.
-The inherent non-stationarity across agents makes value learning unstable (Yu et al., 2022), and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients (Zhang et al., 2024b).
+The inherent non-stationarity across agents makes value learning unstable [49], and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients [51].
 These inaccuracies propagate into the value function, undermining policy quality and stability.
@@
 The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories.
-This integration yields significantly more accurate value-gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
+This integration yields significantly more accurate value‑gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
 Our work makes the following contributions.
-(1) We leverage PINNs to approximate differential value functions and apply them to solve continuous-time multi-agent reinforcement learning problems, which have rarely been explored by previous studies.
-(2) We introduce a novel value-gradient iteration term that dynamically refines the approximations of value gradients during training.
+(1) We leverage PINNs to approximate differential value functions and apply them to solve continuous‑time multi‑agent reinforcement learning problems, which have rarely been explored by previous studies.
+(2) We introduce a novel value‑gradient iteration term that dynamically refines the approximations of value gradients during training.
 This setting improves the computational accuracy of the value gradients, accelerates learning convergence, and leads to highly accurate value approximations, enabling efficient policy learning.
-(3) We create continuous-time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo.
-The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value-gradient learning in high-dimensional multi-agent systems.
+(3) We create continuous‑time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo.
+The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value‑gradient learning in high‑dimensional multi‑agent systems.
 2 Related Work
 2.1 Continuous-Time Reinforcement Learning
-CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings Jia and Zhou (2022); Lee and Sutton (2021); Wallace and Si (2023); Yang et al.
-(2021); Wiltzer et al.
-(2022); Vamvoudakis and Lewis (2010); Bian and Jiang (2021); Faradonbeh and Faradonbeh (2023).
+CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings [18, 23, 40, 47, 46, 38, 4, 11].
 These studies aim to optimize policies in continuous-time domains, which avoids time discretization and achieves higher fidelity in control applications such as robotics and navigation.
-For instance, Bian and Jiang (2021) proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy.
-Similarly, Faradonbeh and Faradonbeh (2023) designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees.
-Lee and Sutton (2021) designs two policy iteration algorithms that compute comprehensive solutions to HJB equations.
-In contrast, Jia and Zhou (2022) introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization.
-Yang et al.
-(2021) develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics.
-Shilova et al.
-(2024) utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings.
-Yildiz et al.
-(2021) integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization.
+For instance, [4] proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy.
+Similarly, [11] designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees.
+[23] designs two policy iteration algorithms that compute comprehensive solutions to HJB equations.
+In contrast, [18] introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization.
+[47] develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics.
+[34] utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings.
+[48] integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization.
 CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings.
-Luviano and Yu (2017) solves multi-agent pathfinding problem using fuzzy Q-learning, while Jiang et al.
-(2023) proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems.
+[26] solves multi-agent pathfinding problem using fuzzy Q-learning, while [19] proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems.
 Beyond these examples, only a limited number of studies have addressed CT-MARL, which highlights the importance of our contributions.
 2.2 Solving HJB Equations via PINNs
-In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations Crandall and Lions (1983), which are the first-order nonlinear parabolic PDEs.
-However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD Osher et al.
-(2004); Osher and Shu (1991).
-Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth Weinan et al.
-(2021).
-PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual Han and Long (2020); Han et al.
-(2018), PDE residual Bansal and Tomlin (2021); Zhang et al.
-(2024b), and supervised data derived from numerical solvers Nakamura-Zimmerer et al.
-(2021); Zhang et al.
-(2024a).
-Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments Mukherjee and Liu (2023).
+In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations [7], which are the first-order nonlinear parabolic PDEs.
+However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD [30, 29].
+Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth [45].
+PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual [14, 15], PDE residual [1, 51], and supervised data derived from numerical solvers [28, 50].
+Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments [27].
 However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
@@
 3.1 Problem Formulation
-In this paper, we focus on multi-agent cooperative settings.
-Following the continuous-time control system framework Doya (2000a); Yildiz et al.
-(2021); Lee and Sutton (2021), we formulate the continuous-time multi-agent problem as a tuple:
+In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework [8, 48, 23], we formulate the continuous-time multi-agent problem as a tuple:
 $\mathcal{M}=\bigl\langle\mathcal{X},\{\mathcal{U}_{i}\}_{i=1}^{N},N,f,r,g,\{t_{k}\}_{k\geq 0},T,\rho\bigr\rangle.$ (1)
@@
 Given $u=(u_{1},\ldots,u_{N})$ as a joint control input, the optimal global value function is defined as:
-$V(x)=\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau)d\tau$ (2)
+$V(x,t)=\max_{u\in\mathcal{U}}\int_{t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau)d\tau+g(x_{T})$ (2)
 3.2 HJB and Policy Learning
@@
 First, we present the following Lemma to show that the optimal value function $V(x,t)$, as defined in Eq.
-(2), is the optimal solution to HJB PDEs Crandall and Lions (1983) in the context of cooperative multi-agent settings.
+(2), is the optimal solution to HJB PDEs [7] in the context of cooperative multi-agent settings.
 Lemma 3.1 (HJB for Multi-agent Systems).
@@
 (3).
-However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 Bui et al.
-(2022).
+However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 [6].
 To address this challenge, we adopt PINNs, which approximate value functions to satisfy both PDE residuals and boundary conditions.
@@
 Lemma 3.2 (Instantaneous Advantage).
-Assume the one-step Q-function over a short interval $\delta t>0$ be
+Assume the one‑step Q‑function over a short interval $\delta t>0$ be
 $Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl(x(t+\delta t),t+\delta t\bigr).$ (6)
@@
 The performance of continuous-time control policies depends critically on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x,t)$.
-Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs Zhang et al.
-(2024b); Mukherjee and Liu (2023); Eberhard et al.
-(2024).
+Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs [51, 27, 10].
 However, the gradients obtained from standard PINN training are often noisy or misaligned with true trajectory behavior.
-Although prior work Zhang et al.
-(2024b) improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings.
+Although prior work [51] improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings.
 To address this, we introduce the VGI module that iteratively refines the value gradient approximations during training.
@@
 3.4.1 Dynamics Model and Reward Model
-In a continuous-time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice.
-Instead, we adopt a discrete-time model-based approach Sutton (1991); Kaiser et al.
-(2019); Hafner et al.
-(2019) that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
+In a continuous‑time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice.
+Instead, we adopt a discrete‑time model‑based approach [36, 20, 13] that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
 $\mathcal{L}_{\text{dyn}}=\bigl\|f_{\psi}(x_{t},u_{t},\Delta t)-x_{t+\Delta t}\bigr\|_{2}^{2}.$ (14)
-After learning $f_{\psi}$, we recover the continuous-time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
+After learning $f_{\psi}$, we recover the continuous‑time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
 Similarly, we fit a reward network $r_{\phi}(x_{t},u_{t},\Delta t)$ to the observed instantaneous reward $r_{t}$:
@@
 $\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{res}}}_{\text{HJB residual}}\;+\;\underbrace{\lambda_{\text{term}}\,\mathcal{L}_{\text{term}}}_{\text{terminal BC}}\;+\;\underbrace{\lambda_{\text{anchor}}\,\mathcal{L}_{\text{anchor}}}_{\text{TD anchor}}\;+\;\underbrace{\lambda_{g}\,\mathcal{L}_{\text{vgi}}}_{\text{VGI consistency}}.$ (18)
-Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value-bootstrap anchoring, and gradient consistency.
+Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value‐bootstrap anchoring, and gradient consistency.
 In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory.
@@
 4 Experimental Results
-We evaluate our Value Iteration via PINN (VIP) method on two continuous-time multi-agent benchmarks: MPE Lowe et al.
-(2017b) and multi-agent MuJoCo Todorov et al.
-(2012); Peng et al.
-(2021).
+We evaluate our Value Iteration via PINN (VIP) method on two continuous‑time multi‑agent benchmarks: MPE [25] and multi‑agent MuJoCo [37, 31].
 In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations.
 This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP.
-Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous-time RL baselines in these environments?
+Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous‑time RL baselines in these environments?
 (2) VGI ablation: How much does the VGI module contribute to final performance and training stability?
@@
 4.1 Baseline Methods
-To evaluate our CT-MARL framework VIP, we compare against three continuous-time policy iteration baselines and include an ablated variant of our method without VGI: CT-MBRL (ODE) Yildiz et al.
-(2021): A continuous-time model-based RL approach that learns system dynamics via Bayesian neural ODEs.
-This method uses an actor-critic framework to approximate state-value functions and derive the continuous-time optimal policies.
+To evaluate our CT‑MARL framework VIP, we compare against three continuous‑time policy iteration baselines and include an ablated variant of our method without VGI: CT‑MBRL (ODE) [48]: A continuous‑time model‑based RL approach that learns system dynamics via Bayesian neural ODEs.
+This method uses an actor‑critic framework to approximate state‑value functions and derive the continuous-time optimal policies.
 Figure 4: Performance across continuous-time MPE settings. The y-axis shows the mean cumulative reward.
-Differential Policy Iteration (DPI) Lee and Sutton (2021): A model-based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations.
-Integral Policy Iteration (IPI) Lee and Sutton (2021): A partially model-free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement.
+Differential Policy Iteration (DPI) [23]: A model‑based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations.
+Integral Policy Iteration (IPI) [23]: A partially model‑free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement.
 In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions.
@@
 The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL.
-As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations Krishnapriyan et al.
-(2021); Zhang et al.
-(2024b).
+As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations [22, 51].
 Figure 5: VIP performance with ReLU and Tanh activation functions in MuJoCo and MPE settings.
@@
 Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments.
-This observation is consistent with the conclusion in the previous studies Zhang et al.
-(2024a), which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
+This observation is consistent with the conclusion in the previous studies [50], which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
 To further demonstrate the importance of VGI, we revisit the didactic example, generate 400 rollouts from sampled initial states, and compute the average value using models with and without VGI.
@@
 The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks.
-This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies Bansal and Tomlin (2021); Zhang et al.
-(2024b).
+This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies [1, 51].
 The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs.
@@
 7, the best accumulated reward performance is achieved when all loss terms are properly balanced.
-Imbalanced weight settings yield stiffness in the training dynamics of PINNs Wang et al.
-(2021), which makes the dominant loss term (with the largest weights) converge faster than the others Wang et al.
-(2022).
+Imbalanced weight settings yield stiffness in the training dynamics of PINNs [42], which makes the dominant loss term (with the largest weights) converge faster than the others [43].
 Figure 8: CT-MARL and DT-MARL performance under different time intervals.
 In our experiments, such an imbalance causes VIP only to satisfy the PDEs or boundary conditions during training, ultimately leading to poor value approximations.
-Lastly, we evaluate the robustness of VIP and MADDPG Lowe et al.
-(2017a) by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts.
+Lastly, we evaluate the robustness of VIP and MADDPG [24] by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts.
 Fig.
@@
 Specifically, we approximate value functions using HJB-based PINNs and introduce VGI to improve value approximations, thus mitigating the adverse impact of inaccurate value approximations on policy learning.
-We validate our VIP across continuous-time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA.
+We validate our VIP across continuous‑time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA.
 Furthermore, we investigate the importance of activation function choice and loss term weighting for VIP performance.
@@
 Many real-world multi-agent scenarios involve competitive interactions, which have not yet been explored within CT-MARL frameworks.
-In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN Zhang et al.
-(2024b).
+In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN [51].
 As part of future work, we aim to extend VIP to explore competitive CT-MARL problems, particularly those involving state constraints.
@@
 B.1 Coupled Oscillator
-We evaluate on a two-agent coupled spring–damper system. Each agent $i\in\{1,2\}$ controls one mass in a pair of identical oscillators with linear coupling. The continuous-time dynamics are
+We evaluate on a two‐agent coupled spring–damper system. Each agent $i\in\{1,2\}$ controls one mass in a pair of identical oscillators with linear coupling. The continuous‐time dynamics are
 $\begin{aligned} \dot{x}_{i}&=v_{i},\\ \dot{v}_{i}&=-\,k\,x_{i}\;-\;b\,v_{i}\;+\;u_{i},\end{aligned}\qquad i=1,2,$ (19)
@@
 $\dot{x}=A\,x+B\,u,$
-we can compute the exact infinite-horizon LQR solution:
+we can compute the exact infinite‑horizon LQR solution:
   1. 1.
@@
   2. 2.
-    Form the optimal state-feedback gain
+    Form the optimal state‑feedback gain
     $K\;=\;R^{-1}B^{\top}P.$
@@
 B.2 Continuous-Time MPE
-We build on the standard MPE of Lowe et al.
-[2017b], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions.
-In the original MPE each control step advances the physics by a fixed time-step $\Delta t_{\rm fixed}=0.1\,$s.
-To evaluate our continuous-time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
+We build on the standard MPE of [25], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions.
+In the original MPE each control step advances the physics by a fixed time‐step $\Delta t_{\rm fixed}=0.1\,$s.
+To evaluate our continuous‐time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
 $\Delta t_{k}\;\sim\;\mathrm{Uniform}(\Delta t_{\min},\,\Delta t_{\max}).$
@@
 B.3 Continuous-Time Multi-Agent MuJoCo
-For high-dimensional control we adapt the discrete-time Multi-Agent MuJoCo suite (e.g.
+For high‑dimensional control we adapt the discrete‐time Multi‑Agent MuJoCo suite (e.g.
 cooperative locomotion, quadruped rendezvous).
-By default MuJoCo uses an internal physics integrator with a base time-step of 0.010.01\,s and repeats each action for $K_{\rm fixed}=5$ frames, yielding an effective control interval $\Delta t_{\rm fixed}=0.05\,$s.
+By default MuJoCo uses an internal physics integrator with a base time‐step of 0.010.01\,s and repeats each action for $K_{\rm fixed}=5$ frames, yielding an effective control interval $\Delta t_{\rm fixed}=0.05\,$s.
 We instead sample the number of frames per control step,
@@
 Across both tasks the Tanh implementation converges faster and attains a higher plateau reward, whereas the ReLU version peaks earlier and then undergoes a mild performance decay.
-The observation aligns with the earlier Tanh-versus-ReLU analysis reported in the main paper Fig.
+The observation aligns with the earlier Tanh‐versus‐ReLU analysis reported in the main paper Fig.
 5: smoother activation functions mitigate gradient saturation and promote more stable policy updates.

### v3

@@
 Existing reinforcement learning (RL) methods struggle with complex dynamical systems that demand interactions at high frequencies or irregular time intervals.
-Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differential value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation.
+Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differentiable value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation.
 While CTRL has shown promise, its applications have been largely limited to the single-agent domain.
-This limitation stems from two key challenges: (i) conventional solution methods for HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with HJB-based learning approaches, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training.
+This limitation stems from two key challenges: (i) conventional methods for solving HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with learning-based approaches to alleviate the CoD, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training.
 In this paper, we propose a CT-MARL framework that uses physics-informed neural networks (PINNs) to approximate HJB-based value functions at scale.
-To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines $\nabla_{x}V$ along trajectories.
-This improves gradient fidelity, in turn yielding more accurate values and stronger policy learning.
+To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines value gradients along trajectories.
+This improves gradient accuracy, in turn yielding more precise value approximations and stronger policy learning.
 We evaluate our method using continuous‑time variants of standard benchmarks, including multi‑agent particle environment (MPE) and multi‑agent MuJoCo.
-Our results demonstrate that our approach consistently outperforms existing continuous‑time RL baselines and scales to complex multi-agent dynamics.
+Our results demonstrate that our approach consistently outperforms existing continuous‑time RL baselines and scales to complex cooperative multi-agent dynamics.
+Code is available at this link
 1 Introduction
-RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation [5], strategy games [39], wireless communications [12, 44] , and traffic coordination [16].
-Most existing RL methods model these interactions in discrete time, where Bellman backup is computed at a fixed time interval.
-However, discrete‑time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading [33] and autonomous driving [21]).
-Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues [41, 34].
-To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time [9, 32].
-However, existing works only focus on single‑agent settings and have not yet been widely explored for multi-agent scenarios.
-In multi-agent domains, each agent must not only interact with the environment but also coordinate with other agents, all while coping with the chllenges like non-stationarity introduced by simultaneously learning policies.
-Within this context, our paper investigates a novel approach to solving CT-MARL problems.
-To better understand the limitations of DTRL and advantages of CTRL for multi-agent scenarios, we present a didactic case as shown in Fig.
-1.
-In this simple continuous‑time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2.
-In contrast, our CT‑MARL algorithm closely follows the ground‑truth trajectory, maintains high returns, and generates accurate control actions for both agents.
+RL has achieved remarkable success in a range of discrete-time single- and multi-agent interaction tasks, including robotic manipulation (Brunke et al., 2022), strategy games (Vinyals et al., 2019), wireless communications (Feriani and Hossain, 2021; Wang et al., 2023), and traffic coordination (Haydari and Yılmaz, 2020).
+However, many real-world domains are inherently continuous-time and operate at high or irregular decision frequencies, such as continuous control (Talbot et al., 2025), autonomous driving (Brechtel et al., 2014), and market trading (Shavandi and Khedmati, 2022).
+Despite this, most existing RL methods formulate decision-making in discrete-time, where Bellman updates are computed at a fixed time interval.
+Such discrete-time RL (DTRL) approximates a continuous-time process by imposing a fixed discretization step, which introduces two inherent limitations.
+First, when the timestep is coarse, the resulting controller becomes non-smooth and leads to suboptimal or unstable behavior (Doya, 2000).
+Second, when the timestep is fine, the number of states and iteration steps become large, which enlarges not only memory storage but also many learning trials (Doya, 2000).
+In addition, as the time step $\Delta t$ approaches to 0, the Bellman operator can become ill-conditioned: the temporal-difference objective functions may be dominated by approximation noise, leading the Bellman updates to become unstable and have a large variance (Wang et al., 2020; Shilova et al., 2024).
+These limitations indicate that performance may critically depend on the chosen discretization method, thereby motivating continuous-time RL (CTRL) methods that avoid timestep-discretization and directly learn value functions in continuous time (Doya, 2000; Rubanova et al., 2019).
 (a) Value Approximation
@@
 (d) Agent 2 Action
-Figure 1: The performance of CT-MARL and DT-MARL is compared on a continuous-time, two-agent coupled oscillator task.
-DT-MARL suffers from significant bias and error when applied in the continuous domain.
+Figure 1: The performance of our CT-MARL and DT-MARL is compared on a continuous-time, two-agent coupled oscillator task.
+In the discrete-time setting, DT-MARL trained with MADDPG can achieve near-optimal performance.
+However, when transferred to the continuous-time domain, MADDPG suffers from significant bias and error, resulting in poor approximations.
 In contrast, CT-MARL yields smoother actions, higher rewards, and more accurate value approximations, closely aligning with the analytical LQR ground truth.
-Unlike Bellman operator-based DTRL [2], CTRL leverages HJB PDEs to compute differential value functions [27, 34].
-However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems [3], especially where the computational complexity grows exponentially with the state dimension in multi-agent systems.
-PINNs have emerged as a powerful tool to circumvent CoD [45], and offer convergence guarantees for problems with smooth solutions [35, 17].
+Unlike Bellman operator-based DTRL (Bellman, 1966), CTRL leverages HJB Partial Differential Equation (PDEs) to compute differential value functions (Mukherjee and Liu, ; Shilova et al., 2024).
+However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or level set method (Osher et al., 2004)) suffers from CoD in high-dimensional dynamical systems (Bellman et al., 1965), especially where the computational complexity grows exponentially with the state dimension in multi-agent systems.
+PINNs have emerged as a powerful tool to circumvent CoD (Weinan et al., 2021), and offer convergence guarantees for problems with smooth solutions (Shin et al., 2020; Ito et al., 2021).
 To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
-However, even with PINNs to alleviate the CoD in CT-MARL, learning accurate centralized value functions under the centralized training decentralized execution (CTDE) framework remains challenging.
-The inherent non-stationarity across agents makes value learning unstable [49], and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients [51].
-These inaccuracies propagate into the value function, undermining policy quality and stability.
-To address this limitation, we propose a novel learning approach that combines PINN and VGI information, which is used to optimize the value learning.
+Yet existing CTRL methods focus almost exclusively on single-agent settings and do not extend naturally to MARL.
+In multi-agent domains, each agent must interact with the environment while simultaneously adapting to other learning agents, introducing severe non-stationarity, making value learning substantially harder (Yu et al., 2022).
+Even with PINNs, accurately learning value functions under the centralized training decentralized execution (CTDE) framework remains challenging: PDE-based residual losses alone often produce biased or inaccurate value gradients, where coupled dynamics amplify approximation errors (Zhang et al., 2024b).
+These inaccuracies propagate through the PDE residuals, degrade value learning, and ultimately deteriorate policy optimization.
+To address this limitation, we propose a novel learning approach that combines PINN and VGI to optimize the value learning.
 The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories.
-This integration yields significantly more accurate value‑gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
+To better understand the limitations of DTRL and advantages of our CTRL algorithms for multi-agent scenarios, we present a didactic case as shown in Fig.
+1.
+In this simple continuous‑time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2.
+In contrast, our CT‑MARL algorithm closely follows the ground‑truth trajectory, maintains high returns, and generates accurate control actions for both agents.
 Our work makes the following contributions.
-(1) We leverage PINNs to approximate differential value functions and apply them to solve continuous‑time multi‑agent reinforcement learning problems, which have rarely been explored by previous studies.
+(1) We leverage PINNs to approximate differential value functions and apply them to solve cooperative continuous‑time multi‑agent reinforcement learning problems within deterministic systems, which have rarely been explored by previous studies.
 (2) We introduce a novel value‑gradient iteration term that dynamically refines the approximations of value gradients during training.
@@
 2.1 Continuous-Time Reinforcement Learning
-CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings [18, 23, 40, 47, 46, 38, 4, 11].
-These studies aim to optimize policies in continuous-time domains, which avoids time discretization and achieves higher fidelity in control applications such as robotics and navigation.
-For instance, [4] proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy.
-Similarly, [11] designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees.
-[23] designs two policy iteration algorithms that compute comprehensive solutions to HJB equations.
-In contrast, [18] introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization.
-[47] develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics.
-[34] utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings.
-[48] integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization.
-CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings.
-[26] solves multi-agent pathfinding problem using fuzzy Q-learning, while [19] proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems.
-Beyond these examples, only a limited number of studies have addressed CT-MARL, which highlights the importance of our contributions.
+CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings.
+These studies aim to optimize policies in continuous-time domains, avoiding time discretization and yielding more accurate control actions in robotics and navigation.
+For instance, Bian and Jiang (2021) proposes a continuous-time value iteration algorithm for solving HJB equations without the need for a stabilizing initial policy, while Faradonbeh and Faradonbeh (2023) introduces an HJB-based actor-critic network with theoretical guarantees for the infinite-horizon case.
+Similarly, Lee and Sutton (2021) develops two policy iteration algorithms that compute comprehensive solutions to HJB equations.
+In contrast, Jia and Zhou (2022a) introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization.
+Building on this line, Jia and Zhou (2022b; 2023) uses rigorous algorithms and strong theoretical foundations to study single-agent CTRL with stochastic dynamics.
+Other approaches include Yang et al.
+(2021), which develops a robust actor–critic framework for nonlinear systems with unmodeled dynamics; Shilova et al.
+(2024), which leverages PINNs with $\epsilon$-scheduling to approximate value functions and empirically outperforms discrete-time RL baselines; and Yildiz et al.
+(2021), which integrates neural ODEs with Bayesian inference to model uncertainty in state evolution and proposes a continuous-time actor–critic algorithm that mitigates challenges such as Q-function vanishing and poor discretization.
+In contrast, CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings.
+For example, Luviano and Yu (2017) solves the multi-agent pathfinding problem using fuzzy Q-learning, while Jiang et al.
+(2023) proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems.
+Beyond these examples, only a limited number of studies have addressed CT-MARL, highlighting the importance of our contributions.
 2.2 Solving HJB Equations via PINNs
-In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations [7], which are the first-order nonlinear parabolic PDEs.
-However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD [30, 29].
-Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth [45].
-PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual [14, 15], PDE residual [1, 51], and supervised data derived from numerical solvers [28, 50].
-Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments [27].
+In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations (Crandall and Lions, 1983), which are the first-order nonlinear parabolic PDEs.
+However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD (Osher et al., 2004; Osher and Shu, 1991).
+Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth (Weinan et al., 2021).
+PINNs approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residuals (Han and Long, 2020; Han et al., 2018), PDE residuals (Bansal and Tomlin, 2021; Zhang et al., 2024a; b), and supervised data derived from numerical solvers (Nakamura-Zimmerer et al., 2021).
+Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments (Mukherjee and Liu, ).
 However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
 3 Methodology
+In this section, we present our Value Iteration via PINN (VIP) framework for solving CT-MARL problems.
+1) We first formulate the problem setting as the cooperative multi-agent deterministic system in continuous-time.
+By adopting a CTDE paradigm, 2) we approximate the HJB-based value function using a PINN equipped with residual and anchor losses.
+To further enhance the stability and accuracy of both the value and its gradient, we incorporate additional VGI loss that iteratively refines both value and value-gradient accuracy along sampled trajectories.
+Building upon this critic, 3) we derive a continuous-time advantage function directly from the HJB residual to guide decentralized policy updates, which enables each agent to update its actor in a manner consistent with continuous-time optimality conditions.
 3.1 Problem Formulation
-In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework [8, 48, 23], we formulate the continuous-time multi-agent problem as a tuple:
-$\mathcal{M}=\bigl\langle\mathcal{X},\{\mathcal{U}_{i}\}_{i=1}^{N},N,f,r,g,\{t_{k}\}_{k\geq 0},T,\rho\bigr\rangle.$ (1)
+In this paper, we focus on multi-agent cooperative settings.
+Following the continuous-time control system framework (Yildiz et al., 2021; Lee and Sutton, 2021), we formulate the continuous-time multi-agent problem as a tuple:
+$\mathcal{M}=\bigl\langle\mathcal{X},\{\mathcal{U}_{i}\}_{i=1}^{N},N,f,r,\{t_{k}\}_{k\geq 0},\rho\bigr\rangle.$ (1)
 where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is the state space and $\mathcal{U}=\mathcal{U}_{1}\times\dots\times\mathcal{U}_{N}\subseteq\mathbb{R}^{m}$ represents the joint action space of $N$ agents.
 The global state and control input are represented by $x\in\mathcal{X}$ and $u\in\mathcal{U}$.
-Agent interactions occur over a fixed time horizon [0, T].
+Agent interactions occur over an infinite time horizon.
 The multi-agent system evolves according to time-invariant nonlinear dynamics defined by $\dot{x}=f(x,u)$, where $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ is the global dynamics function.
-We define $\pi:\mathcal{X}\times[0,T]\rightarrow\mathcal{U}$ as the decentralized joint policy $\pi=(\pi_{1},\dots,\pi_{N})$.
-All agents share a global reward $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ and a global terminal cost $g:\mathcal{X}\rightarrow\mathbb{R}$.
+We define $\pi:\mathcal{X}\rightarrow\mathcal{U}$ as the decentralized joint policy $\pi=(\pi_{1},\dots,\pi_{N})$.
+All agents share a global reward $r:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$.
 $\rho\in(0,1]$ is the discount factor or time-horizon scaling parameter.
 Unlike standard formulations that assume a fixed time step, we consider a strictly increasing sequence of decision times $\{t_{k}\}_{k\geq 0}$ with variable gaps $\tau_{k}=t_{k+1}-t_{k}>0$.
-In this paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f$ is Lipschitz continuous; $r$ and $g$ is Lipschitz continuous and bounded.
+In this paper, we assume that $\mathcal{U}_{i}$ is compact and convex; $f$ is Lipschitz continuous; $r$ is Lipschitz continuous and bounded.
 Definition 1 (Value Function of Multi-agent Systems).
 Given $u=(u_{1},\ldots,u_{N})$ as a joint control input, the optimal global value function is defined as:
-$V(x,t)=\max_{u\in\mathcal{U}}\int_{t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau)d\tau+g(x_{T})$ (2)
+$V(x)=\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau)d\tau$ (2)
 3.2 HJB and Policy Learning
-For CT-MARL problems, we adopt the HJB theory to solve continuous-time problems.
-In this subsection, we explain how the HJB equation is leveraged to solve the CT-MARL problems.
+For CT-MARL problems, we build on HJB PDEs rather than discrete-time Bellman equations (Bellman, 1966), which are ill-suited to continuous-time settings.
+In this subsection, we explain how the HJB equations are leveraged to solve the CT-MARL problems.
 Specifically, we define a value network $V_{\theta}$ parametrized by a set of weights and biases $\theta$, and describe how the global value function $V_{\theta}$ is trained and how each agent’s policy $\pi_{\phi_{i}}$ is updated, such that the overall procedure serves as a continuous-time analogue of actor–critic policy iteration.
-3.2.1 Critic Learning with HJB.
-First, we present the following Lemma to show that the optimal value function $V(x,t)$, as defined in Eq.
-(2), is the optimal solution to HJB PDEs [7] in the context of cooperative multi-agent settings.
+The convergence of HJB-based PINNs for learning $V_{\theta}$ has been established in Meng et al.
+(2024), providing theoretical support for our approach.
+3.2.1 Critic Learning with HJB
+The optimal value function is fully represented by the HJB equation.
+This PDE encodes the fundamental relationship between reward, system dynamics, and value evolution, and serves as the theoretical backbone for our value learning method.
+The following lemma states this connection formally in the multi-agent cooperative setting.
 Lemma 3.1 (HJB for Multi-agent Systems).
-For all $(x,t)\in\mathcal{X}\times[0,T]$, the value function $V(x,t)$ is the optimal solution to satisfy the HJB PDEs and boundary conditions:
-$\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f\bigl(x,u^{*}\bigr)+r\bigl(x,u^{*}\bigr)=0,$ (3)
-where optimal control input $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,t,\nabla_{x}V(x,t))$, and boundary condition is $V(x,T)=g(x)$.
-The Hamiltonian $\mathcal{H}$ is defined as $\mathcal{H}=\nabla_{x}V(x,t)^{\!\top}\!f\bigl(x,u\bigr)+r\bigl(x,u\bigr)$.
+For all $x\in\mathcal{X}$, the value function $V(x)$ is the optimal solution to satisfy the following HJB PDEs:
+$\rho V(x)+\nabla_{x}V(x)^{\!\top}\!f\bigl(x,u^{*}\bigr)+r\bigl(x,u^{*}\bigr)=0,$ (3)
+where optimal control input $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,\nabla_{x}V(x))$.
+The Hamiltonian $\mathcal{H}$ is defined as $\mathcal{H}=\nabla_{x}V(x)^{\!\top}\!f\bigl(x,u\bigr)+r\bigl(x,u\bigr)$.
 The proof is attached in Appendix A.1.
-To approximate differential value functions, we aim to solve the HJ PDEs given in Eq.
-(3).
-However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 [6].
-To address this challenge, we adopt PINNs, which approximate value functions to satisfy both PDE residuals and boundary conditions.
-Specifically, we define HJB PDE residual as
-$\mathcal{R}_{\theta}(x,t)=-\rho V_{\theta}(x,t)+\partial_{t}V_{\theta}(x,t)+\nabla_{x}V_{\theta}(x,t)^{\!\top}f(x,u)+r(x,u).$ (4)
-and minimize both residual loss and boundary condition loss towards zero during model refinement.
-$\mathcal{L}_{\text{res}}=\bigl\|\mathcal{R}_{\theta}(x,t)\bigr\|_{1},\quad\mathcal{L}_{\text{term}}=\bigl\|V_{\theta}(x,T)-g(x)\bigr\|_{2}^{2}.$ (5)
+To approximate differentiable value functions, we solve the HJB PDEs in Eq.
+3.
+Since conventional numerical methods become intractable beyond six state dimensions (Bui et al., 2022), we instead employ PINNs, which approximate value functions by minimizing PDE residuals.
+Specifically, we define the HJB PDE residuals as
+$\mathcal{R}_{\theta}(x_{t})=-\rho V_{\theta}(x_{t})+\nabla_{x_{t}}V_{\theta}(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t}).$ (4)
+and minimize the residual loss $\mathcal{L}_{\text{res}}=\bigl\|\mathcal{R}_{\theta}(x_{t})\bigr\|_{1}$ towards zero during model refinement.
 3.2.2 Policy Learning
-While analytical optimal control laws can be derived in some cases by maximizing the Hamiltonian, such closed-form solutions might not be available in complex multi-agent systems such as MPE or MuJoCo.
+While analytical optimal control laws can be derived in some cases by maximizing the Hamiltonian (Nakamura-Zimmerer et al., 2021; Shilova et al., 2024), such closed-form solutions are not available in complex multi-agent systems like MPE or MuJoCo.
 To overcome this challenge, we use an actor network to generate control inputs, replacing the need for analytical controls in the critic network used to approximate value functions.
 The actor and critic networks are refined iteratively until convergence, enabling the actor network to approximate optimal control policies.
-During training, we compute a continuous-time advantage function derived by residual $\mathcal{R}_{\theta}(x,t)$.
-This advantage function is used for policy gradient, where each agent’s decentralized policy $\pi_{\phi_{i}}(u_{i}\mid x,t)$ is optimized to maximize long-term return.
+During training, we compute a continuous-time advantage function derived by residual $\mathcal{R}_{\theta}(x_{t})$.
+This advantage function is used for policy gradient, where each agent’s decentralized policy $\pi_{\phi_{i}}(u_{i}~|~x_{t})$ is optimized to maximize long-term return.
 Lemma 3.2 (Instantaneous Advantage).
 Assume the one‑step Q‑function over a short interval $\delta t>0$ be
-$Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl(x(t+\delta t),t+\delta t\bigr).$ (6)
+$Q(x_{t},u_{t})=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}V\bigl(x_{t+\delta t}\bigr).$ (5)
 Then the instantaneous advantage satisfies
-$A(x,t,u)=-\rho V(x,t)+\partial V_{t}(x,t)+\nabla_{x}V(x,t)^{\!\top}\,f(x,u)+r(x,u).$ (7)
-The proof is attached at Appendix A.2.
-With the critic’s instantaneous advantage:
-$A_{\theta}(x,t,u)=-\rho V_{\theta}(x,t)+\partial_{t}V_{\theta}(x,t)+\nabla_{x}V_{\theta}(x,t)^{\!\top}f(x,u)+r(x,u),$ (8)
-we update each agent’s policy network $\pi_{\phi_{i}}(u_{i}\mid x,t)$ in a decentralized fashion. For agent $i$, we minimize the negative expected advantage under the joint policy:
-$\mathcal{L}_{p_{i}}=-A_{\theta}\bigl(x,t,u\bigr)\log\pi_{\phi_{i}}\bigl(u_{i}\mid x,t\bigr),$ (9)
+$A(x_{t},u_{t})=-\rho V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}\,f(x_{t},u_{t})+r(x_{t},u_{t}).$ (6)
+The proof is attached in Appendix A.2.
+With the critic’s instantaneous advantage
+$A_{\theta}(x_{t},u_{t})=-\rho V_{\theta}(x_{t})+\nabla_{x_{t}}V_{\theta}(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t}),$ (7)
+we update each agent’s policy network $\pi_{\phi_{i}}(u_{i}~|~x_{t})$ in a decentralized fashion. For agent $i$, we minimize the negative expected advantage under the joint policy
+$\mathcal{L}_{p_{i}}=-A_{\theta}\bigl(x_{t},u_{t}\bigr)\log\pi_{\phi_{i}}\bigl(u_{i}~|~x_{t}\bigr),$ (8)
 Here, $u=(u_{i},u_{-i})$ denotes the joint action, where $u_{i}\sim\pi_{\phi_{i}}$ is sampled from agent $i$’s policy, and $u_{-i}$ represents the actions of all other agents, sampled as $u_{-i}\sim\pi_{\phi_{-i}}$.
+Since our actor is trained by using the advantage function, it is important to ensure that this update direction leads to policy improvement.
+The following lemma formalizes this property and shows that a gradient step on our actor loss yields a value-increasing policy update.
 Lemma 3.3 (Policy Improvement).
 Let $\pi_{\text{old}}$ be the current joint policy and $\pi_{\text{new}}$ the updated policy after one gradient step on the actor loss $\mathcal{L}_{p}$ with sufficiently small step size. Then:
-$Q^{\pi_{\text{new}}}(x,t,u)\geq Q^{\pi_{\text{old}}}(x,t,u).$ (10)
-The proof can be found at Appendix A.3.
+$Q^{\pi_{\text{new}}}(x_{t},u_{t})\geq Q^{\pi_{\text{old}}}(x_{t},u_{t}).$ (9)
+The proof can be found in Appendix A.3.
 3.3 Value Gradient Iteration Module
-The performance of continuous-time control policies depends critically on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x,t)$.
-Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs [51, 27, 10].
-However, the gradients obtained from standard PINN training are often noisy or misaligned with true trajectory behavior.
-Although prior work [51] improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings.
-To address this, we introduce the VGI module that iteratively refines the value gradient approximations during training.
+The performance of continuous-time control policies depends explicitly on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x)$.
+Recent studies have demonstrated that the accuracy of the value directly affects the learned control policies (Zhang et al., 2024b).
+However, enforcing the HJB PDE through residual loss alone does not guarantee accurate value gradients.
+Because in high-dimensional multi-agent systems, small gradient errors can be significantly amplified by coupled dynamics, leading to inaccurate policy gradients.
+To address this limitation, we introduce the VGI module, which explicitly propagates value gradients along sampled trajectories and refines the gradient estimates in a self-consistent manner.
+By iteratively updating $\nabla_{x}V(x)$ using the local dynamics and the current value approximation, VGI provides a trajectory-aligned correction signal that complements the global PDE constraint, resulting in substantially more accurate value-gradient learning.
 Definition 2 (VGI Gradient Estimator).
 Given a small time step $\Delta t$, the VGI estimator of the value gradient at $(x_{t},u_{t})$ is defined by
-$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\!\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$ (11)
+$\nabla_{x_{t}}V(x_{t})=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f(x_{t},u_{t})^{\!\top}\nabla_{x_{t+\Delta t}}V(x_{t+\Delta t}).$ (10)
 The VGI target in Eq.
-(11) can be interpreted as a one-step unrolling of the Bellman equation in the space of gradients.
+10 can be interpreted as a one-step unrolling of the Bellman equations in the space of gradients.
 The first term captures the instantaneous contribution of the local reward gradient, while the second term propagates the downstream value information through the Jacobian of the system dynamics.
@@
 The derivation process is posted at Appendix A.4.
+To justify why the VGI refinement is stable and converges to a promising gradient estimate, we analyze its update rule as a fixed-point iteration.
 Theorem 3.4 (Convergence of VGI).
 Let $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be defined as
-$G(\zeta)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\top}\zeta,$ (12)
-and assume the dynamics $\|\nabla_{x}f(x_{t},u_{t})\|$ is bounded .
+$G(\zeta)=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f(x_{t},u_{t})^{\top}\zeta,$ (11)
+and assume the dynamics $\|\nabla_{x_{t}}f(x_{t},u_{t})\|$ is bounded .
 Then $G$ is a contraction, and the sequence $\zeta^{(k+1)}=G(\zeta^{(k)})$ converges to a unique fixed point $\zeta^{*}\in\mathbb{R}^{d}$.
-The proof is detailed at Appendix A.5
-Rather than introducing a separate network to predict value gradients, we directly compute the automatic derivative of the shared PINN critic $V_{\theta}(x,t)$.
+The proof is detailed in Appendix A.5
+Rather than introducing a separate network to predict value gradients, we directly compute the automatic derivative of the shared PINN critic $V_{\theta}(x_{t})$.
 This gradient is then trained to match the VGI-generated target defined in Eq.
-(11).
+10.
 Specifically, we minimize the mean squared error between the computed and target gradients:
-$\mathcal{L}_{\text{vgi}}=\bigl\|\nabla_{x}V_{\theta}(x_{t},t)-\hat{g}_{t}\bigr\|_{2}^{2},$ (13)
-where $\hat{g}_{t}=\nabla_{x}r_{\phi}(x_{t},u_{t},\Delta t)\Delta t+e^{-\rho\Delta t}\nabla_{x}f_{\psi}(x_{t},u_{t},\Delta t)^{\!\top}\nabla_{x}V_{\theta}(x_{t+\Delta t},t+\Delta t).$ Here, $r_{\phi}(x,u,\Delta t)$ denotes a reward models and $f_{\psi}(x,u,\Delta t)$ represents a dynamics model, where $\phi$ and $\psi$ are respective network parameters.
+$\mathcal{L}_{\text{vgi}}=\bigl\|\nabla_{x_{t}}V_{\theta}(x_{t})-\hat{g}_{t}\bigr\|_{2}^{2},$ (12)
+where $\hat{g}_{t}=\nabla_{x_{t}}r_{\phi}(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f_{\psi}(x_{t},u_{t})^{\!\top}\nabla_{x_{t+\Delta t}}V_{\theta}(x_{t+\Delta t}).$ Here, $r_{\phi}(x_{t},u_{t})$ denotes a reward model and $f_{\psi}(x_{t},u_{t})$ represents a dynamics model, where $\phi$ and $\psi$ are respective network parameters.
 3.4 Implementation Details
@@
 In a continuous‑time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice.
-Instead, we adopt a discrete‑time model‑based approach [36, 20, 13] that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
-$\mathcal{L}_{\text{dyn}}=\bigl\|f_{\psi}(x_{t},u_{t},\Delta t)-x_{t+\Delta t}\bigr\|_{2}^{2}.$ (14)
-After learning $f_{\psi}$, we recover the continuous‑time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
-Similarly, we fit a reward network $r_{\phi}(x_{t},u_{t},\Delta t)$ to the observed instantaneous reward $r_{t}$:
-$\mathcal{L}_{\text{rew}}=\bigl\|r_{\phi}(x_{t},u_{t},\Delta t)-r_{t}\bigr\|_{2}^{2}.$ (15)
+Instead, we adopt a discrete‑time model‑based approach (Sutton, 1991; Hafner et al., 2019) that we train a neural network $f_{\psi}(x_{t},u_{t})$ to predict the next state $x_{t+\Delta t}$ via
+$\mathcal{L}_{\text{dyn}}=\bigl\|f_{\psi}(x_{t},u_{t})-x_{t+\Delta t}\bigr\|_{2}^{2}.$ (13)
+After learning $f_{\psi}$, we recover the continuous‑time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t})-x_{t}}{\Delta t}$.
+Similarly, we fit a reward network $r_{\phi}(x_{t},u_{t})$ to the observed instantaneous reward $r_{t}$:
+$\mathcal{L}_{\text{rew}}=\bigl\|r_{\phi}(x_{t},u_{t})-r_{t}\bigr\|_{2}^{2}.$ (14)
 Both $f_{\psi}$ and $r_{\phi}$ are trained jointly, enabling us to compute the VGI module’s target.
 3.4.2 Anchor Loss for Critic Network
-In addition to the HJB residual and terminal-condition losses, we incorporate a TD-style anchor loss to improve both the stability and accuracy of value learning.
-While the residual loss enforces the correctness of the value gradient, it does not constrain the value of $V(x,t)$.
+In addition to the HJB residual loss, we incorporate a TD-style anchor loss to improve both the stability and accuracy of value learning.
+While the residual loss enforces the correctness of the value gradient, it does not constrain the value of $V(x)$.
 Terminal-condition losses can provide such supervision, but they often rely on access to well-defined terminal targets, which may be unavailable in complex continuous control environments such as MuJoCo.
 In these cases, the anchor loss offers an additional source of value landscape, helping the critic produce reasonable value approximations even when terminal rewards are sparse, delayed, or difficult to specify.
-We define the one-step return as:
-$R_{t}=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V_{\theta}\bigl(x_{t+\Delta t},\,t+\Delta t\bigr).$ (16)
-The anchor loss then enforces the value network to match these returns:
-$\mathcal{L}_{\text{anchor}}=\bigl\|V_{\theta}(x_{t},t)-R_{t}\bigr\|_{2}^{2}.$ (17)
-The overall critic objective combines all four losses:
-$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{res}}}_{\text{HJB residual}}\;+\;\underbrace{\lambda_{\text{term}}\,\mathcal{L}_{\text{term}}}_{\text{terminal BC}}\;+\;\underbrace{\lambda_{\text{anchor}}\,\mathcal{L}_{\text{anchor}}}_{\text{TD anchor}}\;+\;\underbrace{\lambda_{g}\,\mathcal{L}_{\text{vgi}}}_{\text{VGI consistency}}.$ (18)
-Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value‐bootstrap anchoring, and gradient consistency.
+We define the one-step return as
+$R_{t}=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V_{\theta}\bigl(x_{t+\Delta t}\bigr).$ (15)
+The anchor loss then enforces the value network to match these returns
+$\mathcal{L}_{\text{anchor}}=\bigl\|V_{\theta}(x_{t})-R_{t}\bigr\|_{2}^{2}.$ (16)
+The overall critic objective combines all four losses
+$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{res}}}_{\text{HJB residual}}\;+\;\underbrace{\lambda_{\text{anchor}}\,\mathcal{L}_{\text{anchor}}}_{\text{TD anchor}}\;+\;\underbrace{\lambda_{g}\,\mathcal{L}_{\text{vgi}}}_{\text{VGI consistency}}.$ (17)
+Here $\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE residuals, value‐bootstrap anchoring, and gradient consistency.
 In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory.
@@
 4 Experimental Results
-We evaluate our Value Iteration via PINN (VIP) method on two continuous‑time multi‑agent benchmarks: MPE [25] and multi‑agent MuJoCo [37, 31].
+We evaluate our Value Iteration via PINN (VIP) method on two continuous‑time multi‑agent benchmarks: MPE (Lowe et al., 2017b) and multi‑agent MuJoCo (Peng et al., 2021).
 In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations.
 This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP.
-Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous‑time RL baselines in these environments?
+Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperform existing continuous‑time RL baselines in these environments?
 (2) VGI ablation: How much does the VGI module contribute to final performance and training stability?
-(3) PINN design choice: How does activation function choice and loss term weighting in critic network affect the performance of VIP?
+(3) PINN design choice: How does activation function choice and loss term weighting in the critic network affect the performance of VIP?
 (4) Time discretization impact: How well do discrete-time and continuous-time methods perform under arbitrary or unfixed time intervals?
+(5) Comparison between VIP and discrete-time baselines: Could discrete-time baselines perform comparable at continuous-time settings?
 Figure 3: $V$ and $\nabla_{x}V$ contour using VIP w/ VGI and w/o VGI in $d_{1}$-$d_{2}$ frame.
 Benchmarks.
-We evaluate our method against baselines across eight experimental settings.
+We evaluate our method against competitive baselines across eight experimental settings involving up to six agents and 113 state dimensions.
 We extend the MPE framework to a continuous-time formulation by using a variable-step Euler integration scheme, where the time interval $\Delta t$ is sampled from a predefined range at each step.
 Experiments are conducted on the cooperative navigation and predator prey environments.
-Similarly, we adapt MuJoCo to continuous-time settings and evaluate on a suite of multi-agent locomotion tasks, including ant (2 $\times$ 4, 2 $\times$ 4d, 4 $\times$ 2), walker, swimmer, and cheetah.
+Similarly, we adapt MuJoCo to continuous-time settings and evaluate on a suite of multi-agent locomotion tasks, including ant (2 $\times$ 4, 2 $\times$ 4d, 4 $\times$ 2), walker, swimmer, and cheetah (6 $\times$ 1).
 Further implementation details of these benchmarks are provided in Appendix B.2 and B.3.
@@
 4.1 Baseline Methods
-To evaluate our CT‑MARL framework VIP, we compare against three continuous‑time policy iteration baselines and include an ablated variant of our method without VGI: CT‑MBRL (ODE) [48]: A continuous‑time model‑based RL approach that learns system dynamics via Bayesian neural ODEs.
+To evaluate our CT‑MARL framework VIP, we compare against four continuous‑time policy iteration baselines and include an ablated variant of our method without VGI: CT‑MBRL (ODE) (Yildiz et al., 2021): a continuous‑time model‑based RL approach that learns system dynamics via Bayesian neural ODEs.
 This method uses an actor‑critic framework to approximate state‑value functions and derive the continuous-time optimal policies.
 Figure 4: Performance across continuous-time MPE settings. The y-axis shows the mean cumulative reward.
-Differential Policy Iteration (DPI) [23]: A model‑based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations.
-Integral Policy Iteration (IPI) [23]: A partially model‑free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement.
+Differential Policy Iteration (DPI) (Lee and Sutton, 2021): a model‑based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations.
+Integral Policy Iteration (IPI) (Lee and Sutton, 2021): a partially model‑free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement.
+HJBPPO (Mukherjee and Liu, ): a recent method that employs a PINN-based critic to approximate the HJB residual and leverages a standard PPO-style policy optimization scheme to guide agent learning.
 In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions.
 Ablation (w/o VGI): an ablated version of VIP without VGI, which isolates the efficacy of value gradient refinement.
+In addition, to evaluate how VIP compares with DTRL methods, we incorporate three widely used baselines in our experiments: MATD3 (Zhan et al., 2021), MAPPO (Yu et al., 2022), and MADDPG (Lowe et al., 2017a).
 4.2 Results Analysis
@@
 The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL.
-As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations [22, 51].
+As traditional time-dependent HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations (Krishnapriyan et al., 2021).
+This limitation becomes even more severe in the infinite-horizon setting, where the HJB formulation is time-independent and terminal losses are no longer available.
 Figure 5: VIP performance with ReLU and Tanh activation functions in MuJoCo and MPE settings.
 To address this limitation, we incorporate the anchor and VGI loss terms in Eq.
-(18), which capture the landscape of value and its gradient so that PINNs converge to the true values.
+17, which capture the landscape of value and its gradient so that PINNs converge to the true values.
 Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments.
-This observation is consistent with the conclusion in the previous studies [50], which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
+This observation is consistent with the conclusion in the previous studies (Zhang et al., 2024a), which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
 To further demonstrate the importance of VGI, we revisit the didactic example, generate 400 rollouts from sampled initial states, and compute the average value using models with and without VGI.
+Figure 6: Ablation study of different loss terms in critic network for cooperative predator prey.
 As shown in Fig.
@@
 6, which highlights the critical role of PINN in refining value and policy networks.
-Figure 6: Ablation study of different loss terms in critic network for cooperative predator prey.
 Choice of PINN Design.
@@
 Specifically, we train VIP using ReLU and Tanh activations in both MuJoCo and MPE environments and report the accumulated rewards in Fig.
-5.
+5 and Fig.
+11 (Appendix C.2).
 The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks.
-This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies [1, 51].
+This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies (Bansal and Tomlin, 2021; Zhang et al., 2024b).
 The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs.
-PINNs typically use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs.
-The PINN residuals, including PDEs and boundary conditions, are further optimized using gradient descent.
+Figure 7: VIP performance with different weight settings in critic losses.
+PINNs often use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs.
+The PINN residuals, including PDEs, are further optimized using gradient descent.
 Smooth activation functions like Tanh support stable and accurate gradient flow throughout training, enabling more effective value approximations.
-Figure 7: VIP performance with different weight settings in critic losses.
 In contrast, VIP with ReLU often encounters zero-gradient regions during backpropagation, which results in gradient explosion for deeper network architectures or degrades the learning of value functions due to insufficient nonlinearity.
@@
 Specifically, we evaluate three configurations in Eq.
-(18): 1) balanced weights for all loss terms; 2) a large weight for boundary condition while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms.
+17: 1) balanced weights for all loss terms; 2) a large weight for anchor loss while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms.
+Figure 8: CT-MARL and DT-MARL performance under different time intervals.
 As shown in Fig.
 7, the best accumulated reward performance is achieved when all loss terms are properly balanced.
-Imbalanced weight settings yield stiffness in the training dynamics of PINNs [42], which makes the dominant loss term (with the largest weights) converge faster than the others [43].
-Figure 8: CT-MARL and DT-MARL performance under different time intervals.
-In our experiments, such an imbalance causes VIP only to satisfy the PDEs or boundary conditions during training, ultimately leading to poor value approximations.
-Lastly, we evaluate the robustness of VIP and MADDPG [24] by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts.
+Imbalanced weight settings yield stiffness in the training dynamics of PINNs (Wang et al., 2021), which makes the dominant loss term (with the largest weights) converge faster than the others (Wang et al., 2022).
+In our experiments, such an imbalance causes VIP only to satisfy the PDEs residuals or anchor loss during training, ultimately leading to poor value approximations.
+Time discretization impact.
+Lastly, we evaluate the robustness of VIP and a well-trained MADDPG model (Lowe et al., 2017a) by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts.
+Notably, in the original discrete-time training setting with a 4-dimensional state and 1-dimensional action space, MADDPG can already achieve near-optimal performance.
 Fig.
@@
 This result highlights the advantage of VIP in continuous-time multi-agent scenarios.
-5 Conclusion and Limitations
+Figure 9: Comparison between VIP and DTRL baselines.
+DTRL algorithms’ comparison.
+Figure 9 compares VIP with three widely used discrete-time MARL baselines (MATD3, MADDPG, MAPPO) across several multi-agent MuJoCo tasks.
+We observe that all discrete-time methods’ performance degrades substantially in tasks such as Ant and HalfCheetah.
+This degradation is not surprising.
+Discrete-time Bellman backups implicitly assume fixed and sufficiently small time steps, and their value estimates become biased or unstable when the underlying system evolves in continuous-time with arbitrary or environment-dependent time intervals.
+Moreover, large integration steps amplify approximation errors in both the critic update and the target network, leading to inaccurate temporal-difference objective functions and unstable actor gradients, which aligns well with the observation in other works (Tallec et al., 2019; Park et al., 2021; De Asis and Sutton, 2024).
+5 Conclusion
 We propose a novel approach that integrates PINNs into the actor-critic framework to solve CT-MARL problems.
@@
 In summary, our proposed VIP offers a promising approach to solve CT-MARL problems.
-However, we notice the limitation of VIP: the current work only focuses on cooperative CT-MARL tasks.
-Many real-world multi-agent scenarios involve competitive interactions, which have not yet been explored within CT-MARL frameworks.
-In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN [51].
-As part of future work, we aim to extend VIP to explore competitive CT-MARL problems, particularly those involving state constraints.
+Since our current formulation is based on the HJB equation, we focus on cooperative settings; extending VIP to HJI-based formulations that handle competitive or mixed-motive scenarios is a compelling direction for future work.
+Acknowledgments
+This work was supported by the National Science Foundation under Grant No. 2418106.
+Ethics Statement
+This work focuses on decision-making under continuous-time multi-agent systems. All experiments are conducted entirely in simulation and do not involve human subjects or personal data.
 Appendix A Mathematical Proof
@@
 Proof.
-Given all $(x,t)\in\mathcal{X}\times[0,T]$ and a small horizon $\Delta t>0$, we apply a Taylor expansion to the definition of $V(x,t)$ in Eq. (2) to derive the HJB PDE as follows:
-$V(x,t)$ $=\max_{u\in\mathcal{U}}\int_{t}^{T}e^{-\rho(\tau-t)}r(x,u)d\tau+g(x_{T})$
-$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+g(x_{T})$
-$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t}^{T-\Delta t}e^{-\rho(s+\Delta t-t)}r(x(s+\Delta t),u(s+\Delta t))ds+g(x_{T})$
-$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}(\max_{u\in\mathcal{U}}\int_{t}^{T-\Delta t}e^{-\rho(s-t)}r(x(s+\Delta t),u(s+\Delta t))ds+g(x_{T}))$
-$\approx\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}(\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+g(x_{T}))$
-$\approx\max_{u\in\mathcal{U}}r(x,u)\Delta t+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}V(x(t+\Delta t),t+\Delta t)$
-$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+(1-\rho\Delta t+o(\Delta t))\max_{u\in\mathcal{U}}(V(x,t)+\partial_{t}V(x,t)\Delta t+\nabla_{x}V(x,t)^{\!\top}\!f(x,u)\Delta t+o(\Delta t))$
-By canceling out $V(x,t)$ on both sides of the above equality, we obtain that
-$-\rho V(x,t)\Delta t+\partial_{t}V(x,t)\Delta t+\max_{u\in\mathcal{U}}(\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u))\Delta t+o(\Delta t)=0$
+Given all $x\in\mathcal{X}$ with a small horizon $\Delta t>0$, we apply a Taylor expansion to the definition of $V(x)$ in Eq. 2 to derive the HJB PDE as follows:
+$V(x)$ $=\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(\tau-t)}r(x,u)d\tau$
+$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau$
+$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(s+\Delta t-t)}r(x(s+\Delta t),u(s+\Delta t))ds$
+$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(s-t)}r(x(s+\Delta t),u(s+\Delta t))ds$
+$\approx\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau$
+$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}V(x(t+\Delta t))$
+$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+(1-\rho\Delta t+o(\Delta t))\max_{u\in\mathcal{U}}(V(x)+\nabla_{x}V(x)^{\!\top}\!f(x,u)\Delta t+o(\Delta t))$
+By canceling out $V(x)$ on both sides of the above equality, we obtain that
+$-\rho V(x)\Delta t+\max_{u\in\mathcal{U}}(\nabla_{x}V(x)^{\!\top}\!f(x,u)+r(x,u))\Delta t+o(\Delta t)=0$
 Dividing by $\Delta t$ and letting $\Delta t\to 0$, we have that
-$-\rho V(x,t)+\partial_{t}V(x,t)+\max_{u\in\mathcal{U}}(\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u))=0.$
-Therefore, the $V(x,t)$ is the optimal solution to the following HJB PDEs:
-$\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f(x,u^{*})+r(x,u^{*})=0,$
-with boundary condition $V(x,T)=g(x)$.
-Here optimal control input is $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,t,\nabla_{x}V(x,t))$, where $\mathcal{H}$ is the Hamiltonian defined as $\mathcal{H}=\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u)$.
+$-\rho V(x)+\max_{u\in\mathcal{U}}(\nabla_{x}V(x)^{\!\top}\!f(x,u)+r(x,u))=0.$
+Therefore, the $V(x)$ is the optimal solution to the following HJB PDEs:
+$\rho V(x)+\nabla_{x}V(x)^{\!\top}\!f(x,u^{*})+r(x,u^{*})=0,$
+Here optimal control input is $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,\nabla_{x}V(x))$, where $\mathcal{H}$ is the Hamiltonian defined as $\mathcal{H}=\nabla_{x}V(x)^{\!\top}\!f(x,u)+r(x,u)$.
 ∎
@@
 Recall the one-step $Q$-function over a short interval $\delta t>0$, where $u$ is the optimal control input.
-$Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}\,V\bigl(x(t+\delta t),\,t+\delta t\bigr).$
-For small $\delta t$ we have the first-order Taylor expansion in both state and time:
-$V\bigl(x(t+\delta t),\,t+\delta t\bigr)=V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f(x,u)\,\delta t+\partial_{t}V(x,t)\,\delta t+o(\delta t).$
+$Q(x_{t},u_{t})=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}\,V\bigl(x_{t+\delta t}\bigr).$
+For small $\delta t$ we have the first-order Taylor expansion in state:
+$V\bigl(x_{t+\delta t}\bigr)=V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}\!f(x_{t},u_{t})\,\delta t+o(\delta t).$
 Similarly, $e^{-\rho\delta t}=1-\rho\,\delta t+o(\delta t)$.
 Plugging both expansions into $Q(x,t,u)$ gives
-$Q(x,t,u)$ $=r(x,u)\,\delta t+\bigl(1-\rho\,\delta t+o(\delta t)\bigr)\Bigl[V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)\,\delta t+\partial_{t}V(x,t)\,\delta t+o(\delta t)\Bigr]$
-$=r(x,u)\,\delta t+V(x,t)+\bigl[\nabla_{x}V(x,t)^{\!\top}f(x,u)+\partial_{t}V(x,t)-\rho V(x,t)\bigr]\delta t+o(\delta t).$
-Subtract $V(x,t)$ and discard the higher-order term:
-$Q(x,t,u)-V(x,t)=[-\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)+r(x,u)]\delta t+o(\delta t).$
+$Q(x_{t},u_{t})$ $=r(x_{t},u_{t})\,\delta t+\bigl(1-\rho\,\delta t+o(\delta t)\bigr)\Bigl[V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})\,\delta t+o(\delta t)\Bigr]$
+$=r(x_{t},u_{t})\,\delta t+V(x_{t})+\bigl[\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})-\rho V(x_{t})\bigr]\delta t+o(\delta t).$
+Subtract $V(x_{t})$ and discard the higher-order term:
+$Q(x_{t},u_{t})-V(x_{t})=[-\rho V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t})]\delta t+o(\delta t).$
 Dividing by $\delta t$ and letting $\delta t\to 0$ yields the instantaneous advantage density
-$A(x,t,u)=\lim_{\delta t\to 0}\frac{Q(x,t,u)-V(x,t)}{\delta t}=-\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)+r(x,u).$
+$A(x_{t},u_{t})=\lim_{\delta t\to 0}\frac{Q(x_{t},u_{t})-V(x_{t})}{\delta t}=-\rho V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t}).$
 This completes the proof. ∎
@@
 We consider the standard policy improvement step, where the new policy is obtained by maximizing the state-action value function:
-$\pi_{\text{new}}(x,t)=\arg\max_{u\in\mathcal{U}}Q^{\pi_{\text{old}}}(x,t,u),$
+$\pi_{\text{new}}(x_{t})=\arg\max_{u\in\mathcal{U}}Q^{\pi_{\text{old}}}(x_{t},u_{t}),$
 where the one-step Q-function with a small $\delta t>0$ is defined as:
-$Q(x,t,u)$ $=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl(x(t+\delta t),t+\delta t\bigr)$
-$=r(x,u)\,\delta t+e^{-\rho\delta t}\mathbb{E}_{u^{\prime}\sim\pi(\cdot\mid x^{\prime})}\left[Q(x^{\prime},t+\delta t,u^{\prime})\right],$
+$Q(x_{t},u_{t})$ $=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}V\bigl(x_{t+\delta t}\bigr)$
+$=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}\mathbb{E}_{u^{\prime}\sim\pi(\cdot\mid x^{\prime})}\left[Q(x_{t+\delta t}^{\prime},u_{t+\delta t}^{\prime})\right],$
 Since we assume the goal state function stays invariant, we only define the $Q^{\pi}=\int_{0}^{T}e^{-\rho t}r_{t}dt$.
 Because the new policy $\pi^{\text{new}}$ yields equal or higher value in expectation:
-$\mathbb{E}_{u\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x,t,u)\right]\geq\mathbb{E}_{u\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x,t,u)\right].$
+$\mathbb{E}_{u\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{t},u_{t})\right]\geq\mathbb{E}_{u\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t},u_{t})\right].$
 Then we can have:
-$\begin{aligned} Q^{\pi_{\text{old}}}&=r_{0}\,\delta t+e^{-\rho\delta t}(\mathbb{E}_{u_{1}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{1},t_{1},u_{1})\right])\\ &\leq r_{0}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{1}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{1},t_{1},u_{1})\right]\right)\\ &=r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{2},t_{2},u_{2})\right]\right)\\ &\leq r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{2},t_{2},u_{2})\right]\right)\\ &=r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho t_{2}}r_{2}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{3}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{3},t_{3},u_{3})\right]\right)\\ &\vdots\\ &\leq\int_{0}^{T}e^{-\rho t}r_{t}dt\\ &=Q^{\pi_{\text{new}}}.\end{aligned}.$
+$\begin{aligned} Q^{\pi_{\text{old}}}&=r_{t_{0}}\,\delta t+e^{-\rho\delta t}(\mathbb{E}_{u_{t_{1}}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t_{1}},u_{t_{1}})\right])\\ &\leq r_{t_{0}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{t_{1}}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{t_{1}},u_{t_{1}})\right]\right)\\ &=r_{t_{0}}\,\delta t+e^{-\rho t_{1}}r_{t_{1}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t_{2}},u_{t_{2}})\right]\right)\\ &\leq r_{t_{0}}\,\delta t+e^{-\rho t_{1}}r_{t_{1}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{t_{2}}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{t_{2}},u_{t_{2}})\right]\right)\\ &=r_{t_{0}}\,\delta t+e^{-\rho t_{1}}r_{t_{1}}\,\delta t+e^{-\rho t_{2}}r_{t_{2}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{t_{3}}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t_{3}},u_{t_{3}})\right]\right)\\ &\vdots\\ &\leq\int_{t}^{\infty}e^{-\rho t}r_{t}dt\\ &=Q^{\pi_{\text{new}}}.\end{aligned}.$
 A.4 Derivation of Definition 2
 Proof.
-We consider the value function defined in Eq. (2) and follow the Proof of Lemma 3.1 to write out the dynamic programming principle of $V(x,t)$ as:
-$V(x_{t},t)=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V(x_{t+\Delta t},t+\Delta t).$
-where $u$ is the optimal control input. Taking the gradient with respect to $x$ on both sides using the chain rule:
-$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$
+We consider the value function defined in Eq. 2 and follow the Proof of Lemma 3.1 to write out the dynamic programming principle of $V(x_{t})$ as:
+$V(x_{t})=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V(x_{t+\Delta t}).$
+where $u_{t}$ is the optimal control input. Taking the gradient with respect to $x$ on both sides using the chain rule:
+$\nabla_{x_{t}}V(x_{t})=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f(x_{t},u_{t})^{\top}\nabla_{x_{t+\Delta t}}V(x_{t+\Delta t}).$
 which matches the estimator proposed in Definition 2. ∎
@@
 Proof.
-From VGI in definition 2, with an Euler step gives $x_{t+\Delta t}=x_{t}+f(x_{t},u_{t})\Delta t+o(\Delta t)$, we have the first-order-accurate VGI approximation:
-$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\bigl[I+\nabla_{x}f(x_{t},u_{t})\Delta t\bigr]^{\!\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$
-Rewrite VGI approximation as the affine map $\zeta=G(\zeta)$ with
-$b=\nabla_{x}r(x_{t},u_{t})\Delta t,\;A=e^{-\rho\Delta t}\bigl[I+\nabla_{x}f(x_{t},u_{t})\Delta t\bigr]^{\!\top},\;\zeta=\nabla_{x}V(\cdot,\cdot).$
-Assume the dynamics have a bounded Jacobian $\|\nabla_{x}f(x_{t},u_{t})\|\leq L_{f}$. Then
+From the definition of VGI in Definition 2, applying a first-order Euler step gives
+$x_{t+\Delta t}=x_{t}+f(x_{t},u_{t})\Delta t+o(\Delta t).$
+This yields the first-order VGI approximation
+$\nabla_{x_{t}}V(x_{t})=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\bigl[I+\nabla_{x_{t}}f(x_{t},u_{t})\Delta t\bigr]^{\!\top}\nabla_{x_{t+\Delta t}}V(x_{t+\Delta t}).$
+We can rewrite this approximation as an affine map
+$\zeta=G(\zeta)$
+where
+$b=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t,\;A=e^{-\rho\Delta t}\bigl[I+\nabla_{x_{t}}f(x_{t},u_{t})\Delta t\bigr]^{\!\top},\;\zeta=\nabla_{x}V(x).$
+In Section 3.1, the joint dynamics function $f(x,u)$ is assumed time-invariant, which implies its Jacobian $\nabla_{x}f$ is also time-independent.
+Consequently, the matrix $A$ is time-invariant, and the update map can be written as a single contraction map
+$G(\zeta)=b+\mathcal{A}\zeta,$
+rather than a family of time-indexed maps $G_{t}$}. Assuming the dynamics have a bounded Jacobian,
+$\|\nabla_{x}f(x_{t},u_{t})\|\leq L_{f},$
+we obtain the bound
 $\|A\|\leq e^{-\rho\Delta t}(1+L_{f}\Delta t)=\beta.$
@@
 7:   for each agent $i=1,\dots,N$ do
-8:     $u_{i}\sim\pi_{\omega_{i}}(u_{i}\mid x,t)$
+8:     $u_{i}\sim\pi_{\omega_{i}}(u_{i}\mid x)$
 9:   end for
@@
 11:   $(x^{\prime},r)\leftarrow\mathrm{env.step}(u)$
-12:   append $(x,t,u,r,x^{\prime})$ to local rollout $\mathcal{R}$
+12:   append $(x,u,r,x^{\prime})$ to local rollout $\mathcal{R}$
 13:   $x\leftarrow x^{\prime}$
@@
 15:  $\triangleright$ Dynamics and Reward Model learning on $\mathcal{R}$
-16:  update $\psi,\phi$ as per the Equation (14) and (15).
+16:  update $\psi,\phi$ as per the Eq. 13 and 14.
 17:  $\triangleright$ Critic update on $\mathcal{R}$
-18:  compute all losses $\mathcal{L}_{\mathrm{res}},\mathcal{L}_{\mathrm{term}},\mathcal{L}_{\mathrm{anchor}},\mathcal{L}_{\mathrm{vgi}}$ by Equation (5), (17), (13).
+18:  compute all losses $\mathcal{L}_{\mathrm{res}},\mathcal{L}_{\mathrm{anchor}},\mathcal{L}_{\mathrm{vgi}}$ by Eq. 17.
 19:  $\theta\leftarrow\theta-\alpha_{V}\nabla_{\theta}(\ldots)$
@@
 21:  for $i=1,\dots,N$ do
-22:   compute $A(x,t,u)$ for all $(x,t,u)\in\mathcal{R}$
-23:   $\omega_{i}\leftarrow\omega_{i}-\alpha_{\pi}\,\nabla_{\omega_{i}}\bigl(-\mathbb{E}_{(x,t,u)\in\mathcal{R}}[\,A(x,t,u)\,\log\pi_{\omega_{i}}(u_{i}\mid x,t)\,]\bigr)$ by Equation (9).
+22:   compute $A(x,u)$ for all $(x,u)\in\mathcal{R}$
+23:   $\omega_{i}\leftarrow\omega_{i}-\alpha_{\pi}\,\nabla_{\omega_{i}}\bigl(-\mathbb{E}_{(x,u)\in\mathcal{R}}[\,A(x,u)\,\log\pi_{\omega_{i}}(u_{i}\mid x)\,]\bigr)$ by Eq. 8.
 24:  end for
@@
 B.2 Continuous-Time MPE
-We build on the standard MPE of [25], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions.
+We build on the standard MPE of Lowe et al.
+(2017b), which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions.
 In the original MPE each control step advances the physics by a fixed time‐step $\Delta t_{\rm fixed}=0.1\,$s.
@@
 The additional evidence from the two Ant variants therefore reinforces our previous claim that Tanh is better suited for value–gradient propagation in VIP.
-Appendix D Hyper-parameters
+C.3 Stochastic Noise &\& Model Mismatch
+Figure 12: Performance under Different Noise Levels.
+Figure 12 evaluates the robustness of VIP under different levels of stochastic perturbation injected directly into the dynamics, according to
+$x_{t+\Delta t}\;=\;f(x_{t},u_{t})\,\Delta t\;+\;\varepsilon_{t},\quad\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}I),$
+where we consider three noise scales: $\sigma^{2}=0.1$ (Low), $0.5$ (Mid), and $1.0$ (High).
+With no or low noise, VIP maintains strong performance across both Predator–Prey and Cooperative Navigation, indicating that the learned continuous-time value function and its gradients remain stable under mild stochasticity.
+As the noise level increases, performance degrades noticeably, which is expected: our method is designed for deterministic continuous-time systems, and stochastic settings fundamentally alter the associated HJB equations (leading to stochastic HJB or HJB–Bellman–Fokker–Planck forms).
+Since VIP does not incorporate additional components for modeling diffusion terms or uncertainty, the degradation at high noise levels is consistent with the theoretical scope of our formulation.
+C.4 Sensitivity Check for $\Delta t$
+Figure 13: Average Return under Different $\Delta t$ in Coupled Oscillator .
+[…]
 Table 1: Hyperparameter settings used in all experiments.
@@
 Random seed 111-120
+Appendix D Hyper-parameters
 As Table LABEL:tab:hp shows, the exploration steps are used to delay the decay of the exploration rate: during the first 1000 steps, the exploration schedule remains fixed to encourage initial exploration.
@@
 Network Input Dimension Architecture and Activation
-Value Network State + Time ($d+1$) FC(128) → FC(128) → FC(1), ReLU or Tanh
-Dynamics Network State + Joint Action + Time ($d+na+1$) FC(128) → FC(128) → FC($d$), ReLU
-Reward Network State + Joint Action + Time ($d+na+1$) FC(128) → FC(128) → FC(1), ReLU
-PolicyNet Observation + Time + Time Interval ($o+2$) FC(128) → FC(128) → FC(64) → FC($a$), ReLU
+Value Network State ($d$) FC(128) → FC(128) → FC(1), ReLU or Tanh
+Dynamics Network State + Joint Action ($d+na$) FC(128) → FC(128) → FC($d$), ReLU
+Reward Network State + Joint Action ($d+na$) FC(128) → FC(128) → FC(1), ReLU
+PolicyNet Observation + Time Interval ($o+1$) FC(128) → FC(128) → FC(64) → FC($a$), ReLU
 Table 2 summarizes the architectures of the neural networks used in our VIP framework.
 All networks are implemented as FC layers with hidden size 128 unless otherwise specified.
-The value network takes the concatenation of the global state and time $(d+1)$ as input and outputs a scalar value.
+The value network takes the concatenation of the global state and time $(d)$ as input and outputs a scalar value.
 In our implementation, we use the Tanh activation function for the value network, as it provides smoother and more stable gradient propagation, which is critical for PINN-based value approximation.
 To validate this choice, we conducted an ablation study comparing Tanh and ReLU activations in the previous section.
+Appendix E The Use of Large Language Models (LLMs)
+We use LLMs as a writing assistant to polish/revise the paper.